### v1

# Prompt Optimization via Retrieved Reasoning Assets and Multi-Agent Analysis

Abstract.
Prompt optimization has emerged as an effective alternative to retraining for improving the performance of Large Language Models (LLMs). However, most existing approaches treat evaluation as a black box, relying solely on numerical scores while offering limited insight into why a prompt succeeds or fails. They also depend heavily on trial-and-error refinements, which are difficult to interpret and control. In this paper, we introduce MA-SAPO, a Multi-Agent framework for Score-Aware Prompt Optimization. Compared to prior methods, MA-SAPO explicitly couples evaluation outcomes with structured reasoning to guide systematic edits. The framework specifically consists of two stages: during the Reasoning Phase, agents collaboratively explain metric scores, diagnose weaknesses, and synthesize targeted refinements that are stored as reusable reasoning assets; during the Test Phase, agents retrieve these assets to analyze optimized prompts and apply only evidence-grounded edits. By turning evaluation signals into interpretable reasoning chains, MA-SAPO produces prompt refinements that are more transparent, auditable, and controllable. Experiments on the HelpSteer1/2 benchmarks demonstrate consistent improvements over single-pass prompting, retrieval-augmented baselines, and prior multi-agent strategies, validating the effectiveness of our approach. [^1] Code is available at: https://anonymous.4open.science/r/MA-SAPO-F313.
Automatic Prompt Optimization, Score-Aware Prompt Optimization, Retrieval-Augmented Generation, Large Language Models (LLMs), Multi-Agent System
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Multi-agent systemsccs: Information systems Query representationccs: Computing methodologies Natural language generationccs: Information systems Relevance assessment
Figure 1. Comparative overview of prompt optimization methods. MA-SAPO enhances interpretability by linking evaluation scores to reasoning-driven and evidence-based refinements.
1. Introduction
Large Language Models (LLMs) have emerged as powerful tools capable of tackling a wide range of tasks, from reasoning and summarization to complex dialogue and code generation (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023). However, their performance remains highly sensitive to the wording, structure, and examples embedded within prompts (Pryzant et al., 2023; He et al., 2024; Errica et al., 2024; Cao et al., 2024; Razavi et al., 2025). Consequently, prompt optimization has rapidly emerged as a practical alternative for improving model behavior, as it avoids the need for costly retraining or parameter updates.
Early research explored single-pass prompt optimization strategies that treated the model itself as the optimizer. For instance, methods such as Chain-of-Thought (CoT) prompting (Wei et al., 2022), role assignment through carefully designed system instructions (Wu et al., 2023; Kong et al., 2023), or structured variants such as Tree-of-Thought (ToT) (Yao et al., 2023), Graph-of-Thought (GoT) (Besta et al., 2024), and step-back prompting (Zheng et al., 2023) demonstrated that reasoning patterns or role signals embedded into prompts could substantially improve downstream performance. While effective, these approaches typically operate in isolation: a single model instance refines prompts without incorporating systematic feedback from prior attempts or external resources.
As LLMs have advanced, multi-agent frameworks have gained traction as a way to diversify perspectives and coordinate multiple roles in the optimization process (Zhou et al., 2025; Zhang et al., 2025; Han et al., 2025). By assigning diverse specialized agents to different tasks such as decomposition, evaluation, or refinement, these systems move beyond single-pass rewriting to more structured optimization pipelines. At the same time, optimization-theoretic methods (Li et al., 2025b; Cui et al., 2025) and label-efficient pipelines (Xiang et al., 2025; Dong et al., 2025) have sought to reduce the reliance on costly feedback while ensuring scalability across tasks.
Despite these advances, several common limitations remain: (1) most existing frameworks reduce evaluation to outcome scores, treating it as a black box and leaving practitioners uncertain about why certain prompts succeed or fail; (2) optimization often relies on repetitive trial-and-error refinements that consume significant computation while offering little transparency; (3) even when reasoning is incorporated, it usually remains implicit and is not transformed into explicit, auditable artifacts that can guide systematic edits; (4) current pipelines largely optimize for higher outcome scores but offer limited interpretability and controllability, preventing users from understanding trade-offs or applying targeted adjustments.
To address these limitations, we propose MA-SAPO, a sequential reasoning framework for prompt optimization that explicitly links metric outcomes to actionable edits through structured reasoning artifacts. Specifically, MA-SAPO operates in two complementary phases. In the reasoning phase, three agents transform annotated scores into progressively enriched artifacts: (1) a Metric Explainer Agent interprets evaluation dimensions, (2) a Diagnostician Agent uncovers error sources and trade-offs, and (3) an Action Synthesizer Agent produces concrete edit directives. These artifacts are stored as reusable reasoning assets. In the test phase, a retrieval-augmented pipeline leverages these assets: (1) an Analyzer Agent contrasts the current prompt with retrieved exemplars to identify improvement opportunities, while (2) a Refiner Agent applies targeted edits supported by diagnostic evidence, ensuring that optimizations are interpretable and controllable.
Extensive experiments on the open-source HelpSteer1/2 benchmarks (Wang et al., 2023, 2024a) demonstrate that MA-SAPO consistently outperforms single-pass prompting, retrieval-augmented generation, and prior multi-agent baselines, highlighting the importance of structured reasoning as a bridge between evaluation and optimization. Our main contributions include:
  • We propose a score-aware multi-agent training pipeline that distills evaluation outcomes into reusable, semi-structured reasoning assets, enabling transparent prompt explanations, fine-grained diagnoses, and actionable edit directives.
  • We design MA-SAPO, a modular and computationally efficient framework that achieves interpretable, auditable, and controllable improvements across diverse datasets and model backbones, while substantially reducing token and API call budgets.
  • We conduct extensive experiments on HelpSteer1 and HelpSteer2 benchmarks, demonstrating consistent gains across multiple evaluation metrics and outperforming six state-of-the-art baselines in both effectiveness and efficiency.
2. Related Work
Single-Pass Prompt Optimization. Prompt optimization has rapidly evolved as a practical means of improving the performance of LLMs without retraining (Wei et al., 2022; Pryzant et al., 2023; Sun et al., 2023; Li et al., 2025a). Early studies highlighted the centrality of prompts as control variables for steering model behavior. Specifically, few-shot prompting (Brown et al., 2020) demonstrated that models could adapt to unseen tasks using in-context examples alone, while retrieval-augmented generation (RAG) (Lewis et al., 2020; Lee et al., 2025) injected external knowledge to improve factuality and adaptability. These two paradigms, context construction and evidence injection, laid the foundation for single-pass prompt optimization methods. Subsequent research explored reasoning-augmented prompting. Chain-of-Thought (CoT) (Wei et al., 2022) introduced intermediate reasoning steps. Building on this idea, Tree-of-Thought (ToT) (Yao et al., 2023) generalized linear reasoning chains into branching structures with self-evaluation via lookahead and backtracking. Other structured strategies, such as Skeleton-of-Thought (SoT) (Ning et al., 2023), separate planning from execution to parallelize generation, while Graph-of-Thought (GoT) (Besta et al., 2024) organizes reasoning trajectories into graph topologies. Evolutionary and heuristic single-pass approaches have also been explored. For instance, PromptBreeder (Fernando et al., 2023) evolves a population of prompts through mutation and selection, while PromptWizard (Agarwal et al., 2024) refines prompts via feedback-driven critique-and-synthesis loops.
Multi-Agent Based Prompt Optimization. Beyond single-pass methods, multi-agent systems have emerged as a powerful paradigm for prompt optimization (Du et al., 2023; Liu et al., 2024; Chen et al., 2024; Li et al., 2024). Early agentic frameworks such as Self-Ask (Press et al., 2022) introduced decomposition by generating sub-questions and integrating external retrieval, laying the groundwork for cooperative prompting strategies. Also, Multi-Agent Debate (MAD) (Du et al., 2023) advanced this idea by allowing agents to iteratively present and critique answers in a roundtable format, thereby improving factual grounding and reasoning quality. Reflection-based frameworks (Shinn et al., 2023; Khan et al., 2024) further enhanced adaptability by storing feedback as memory for subsequent optimization attempts. Building on this foundation, recent systems formalized agent roles and workflows to optimize prompts more systematically (Li et al., 2023). Subsequent work explicitly targeted prompt optimization as the objective. For instance, MASS (Zhou et al., 2025) searches over agent topologies while performing block-level prompt refinement, followed by global optimization. Additionally, MARS (Zhang et al., 2025) leverages a Planner and Teacher–Critic–Student dialogue to iteratively propose, critique, and revise prompts in structured rounds, and also MAPGD (Han et al., 2025) integrates pseudo-gradient feedback with bandit-style exploration, offering sample-efficient optimization through agent collaboration. However, most existing frameworks treat evaluation as a black box with limited interpretability and rely on costly trial-and-error refinements. MA-SAPO addresses this by grounding optimization in score-aware reasoning assets that link evaluation outcomes to actionable, interpretable, and auditable refinements.
Figure 2. Overview of the MA-SAPO framework. The training phase constructs reasoning assets $(\mathcal{R}_{i}=(\mathcal{C}_{i},\mathcal{D}_{i},\mathcal{E}_{i}))$ from annotated prompt–response pairs, while the test phase retrieves top-k training examples and their reasoning assets to refine a new test prompt.
3. Methodology
3.1. Preliminaries
We denote the training dataset as
$\mathcal{D}_{train}=\{\tau_{i}\}_{i=1}^{N},\quad\tau_{i}=(p_{i},r_{i},\mathcal{S}_{i}),$
where each entry $\tau_{i}$ consists of a prompt $p_{i}$, its associated response $r_{i}$, and a set of scores $\mathcal{S}_{i}=\{s_{help},s_{corr},s_{coh},s_{comp},s_{verb}\}$, corresponding to the dimensions of helpfulness, correctness, coherence, complexity, and verbosity. The test set is denoted as
$\mathcal{D}_{test}=\{p_{j}\}_{j=1}^{M},$
where only prompts are available, and corresponding responses are generated during evaluation. Our objective is to generate structured reasoning assets from $\mathcal{D}_{train}$ that can later be retrieved and utilized to optimize new prompts and their responses. Thus, the training phase serves as data generation, while the test phase operates as a retrieval-augmented optimization pipeline.
Throughout this framework, we employ several specialized agents denoted as $G$. Each agent is designed with a distinct functional role: $G_{\text{exp}}$ explains metric-level outcomes, $G_{\text{diag}}$ diagnoses weaknesses and trade-offs, $G_{\text{syn}}$ synthesizes actionable directives, $G_{\text{ana}}$ analyzes retrieved examples, and $G_{\text{ref}}$ refines prompts into optimized versions. These agents are executed in sequential order within their respective phases, and their outputs are explicitly stored for later retrieval and reasoning. The overview of the MA-SAPO framework is shown in Figure 2.
3.2. Training Phase: Reasoning Asset Construction
In the training phase, three agents are executed sequentially for each $\tau_{i}\in\mathcal{D}_{train}$. All three agents operate on the same triplet $\tau_{i}=(p_{i},r_{i},\mathcal{S}_{i})$, ensuring that their reasoning is consistently grounded in both the input and output of the original instance.
3.2.1. Metric Explainer Agent $G_{\text{exp}}$
The Metric Explainer Agent provides natural language justifications for the given scores. Formally, given $\tau_{i}$, the agent produces a reasoning card $\mathcal{C}_{i}$:
(1) $\mathcal{C}_{i}=G_{\text{exp}}(\tau_{i},P_{\text{exp}}),$
where $P_{\text{exp}}$ is the system prompt guiding explanation. Each card explicitly outlines the reason why $\tau_{i}$ received its annotated scores. The reasoning card $\mathcal{C}_{i}$ then serves as input to the next agent, together with $\tau_{i}$.
3.2.2. Diagnostician Agent $G_{\text{diag}}$
The Diagnostician Agent operates on $\tau_{i}$ and the reasoning card $\mathcal{C}_{i}$. It extends $\mathcal{C}_{i}$ by analyzing metric-level weaknesses and trade-offs. $G_{\text{diag}}$ produces a diagnostic summary $\mathcal{D}_{i}$ that identifies: (1) the key causes of low-scoring dimensions, and (2) trade-offs across metrics (e.g., verbosity vs. coherence).
(2) $\mathcal{D}_{i}=G_{\text{diag}}(\tau_{i},\mathcal{C}_{i},P_{\text{diag}}).$
The diagnostic summary $\mathcal{D}_{i}$ is then passed to the next stage, along with $\tau_{i}$.
3.2.3. Action Synthesizer Agent $G_{\text{syn}}$
The Action Synthesizer Agent receives $\tau_{i}$ together with $\mathcal{C}_{i}$ and $\mathcal{D}_{i}$. It converts these enriched insights into actionable edit directives (EDs). Each directive corresponds to a concrete modification strategy for improving $\tau_{i}$:
(3) $\mathcal{E}_{i}=G_{\text{syn}}(\tau_{i},\mathcal{C}_{i},\mathcal{D}_{i},P_{\text{syn}}),$
where $\mathcal{E}_{i}=\{e_{1},e_{2},\dots,e_{m}\}$ is a set of recommended edits.
Definition (Reasoning Assets).
For notational convenience, we define the collection of training outputs for instance $i$ as:
$\mathcal{R}_{i}=(\mathcal{C}_{i},\mathcal{D}_{i},\mathcal{E}_{i}).$
As a result, $\mathcal{R}_{i}$ is stored as the reasoning assets aligned with $\tau_{i}$, forming the retrieval corpus for the test phase. Importantly, these reasoning assets are designed as semi-structured text, which makes them machine-parseable and directly usable by downstream agents in the test phase.
3.3. Test Phase: Retrieval and Reformulation
In the test phase, optimization is performed via retrieval-augmented generation. Given a new prompt $p_{\text{test}}$, we adopt a sparse lexical retriever that ranks training prompts based on term-frequency and inverse document-frequency statistics. Each training prompt $p_{j}$ is represented in the sparse lexical space, and a relevance score is computed between $p_{\text{test}}$ and $p_{j}$ according to token overlap weighted by discriminative importance. The top-$k$ most relevant prompt-response pairs are then retrieved along with their corresponding reasoning assets:
$\{(\tau_{j},\mathcal{R}_{j})\}_{j=1}^{k}.$
3.3.1. Analyzer Agent $G_{\text{ana}}$
The Analyzer Agent compares $p_{\text{test}}$ against retrieved examples to identify improvement points. It outputs an improvement report $\mathcal{A}$:
(4) $\mathcal{A}=G_{\text{ana}}(p_{\text{test}},\{\tau_{i},\mathcal{R}_{j}\}_{j=1}^{k},P_{ana}).$
The Analyzer Agent transforms raw retrieval into structured insights rather than merely transferring edits from similar prompts. The agent highlights concrete weaknesses and improvement opportunities by contrasting the test prompt with retrieved reasoning assets.
3.3.2. Refiner Agent $G_{\text{ref}}$
The Refiner Agent regenerates an optimized prompt $\hat{p}$ by incorporating the improvement report:
(5) $\hat{p}=G_{\text{ref}}(p_{\text{test}},\mathcal{A},P_{\text{ref}}).$
It operationalizes the analyzer’s insights into a concrete optimized prompt. Compared to a direct rewriting approach, which may introduce irrelevant or unjustified changes, the refiner explicitly conditions on the improvement report, producing refinements that are focused, justifiable, and consistent with diagnostic evidence. The final optimized prompt $\hat{p}$ is then used to generate a corresponding optimized response $\hat{r}$.
3.4. Evaluation
[…]
(6) $\text{Score}(\hat{p},\hat{r})=\langle s_{help},s_{corr},s_{coh},s_{comp},s_{verb}\rangle,$
where each $s_{m}\in[0,4]$ for $m\in\mathcal{M}$. These scores are normalized into the $[0,1]$ range and then aggregated into a single composite value. [^2] This evaluation methodology is implemented in our setup in Section 4.2. Additionally, We complement automatic metrics with a human evaluation (Section 5.3) that tests (H1) the usefulness, accuracy, and consistency of the reasoning and (H2) whether optimized prompts preserve the original intent; 14 expert annotators, all data scientists or AI engineers, participated.
4. Experimental Design and Setup
4.1. Datasets
We conduct our experiments on the HelpSteer family of datasets (Wang et al., 2023, 2024a), which provide human-annotated prompt-response pairs for evaluating prompt optimization in Large Language Models (LLMs). Both datasets share the same annotation schema across five quality dimensions, but differ in size and design.
We specifically build the retrieval corpus from the HelpSteer2 training dataset, where each instance $\tau_{i}=(p_{i},r_{i},\mathcal{S}_{i})$ is augmented with the reasoning assets $\mathcal{R}_{i}=(\mathcal{C}_{i},\mathcal{D}_{i},\mathcal{E}_{i})$ produced during our training pipeline. Evaluation is conducted on both HelpSteer1 and HelpSteer2 validation prompts.
HelpSteer1 contains 35.3k training and .79k validation samples, each comprising a prompt, a response, and five human-annotated attributes scored on a 0-4 scale. It served as the first large-scale benchmark for aligning model outputs with human feedback and provides broad coverage of prompt-response styles.
HelpSteer2 is a newer dataset containing 20.3k training and 1.04k validation samples, with higher-quality annotations, multi-turn prompts (about 29% of cases), and additional preference annotations between responses (Wang et al., 2024a). We adopt it as our retrieval corpus since (i) its richer annotations and multi-turn coverage yield more representative reasoning assets, (ii) preference signals provide finer-grained evidence for our Analyzer and Refiner agents, and (iii) it is the current benchmark standard for training and evaluating reward models, ensuring alignment with best practices in the field.
Table 1. Overview of HelpSteer1/2 Annotation Dimensions. Each scored on a 0–4 scale.
Attribute Description
Helpfulness Overall helpfulness of the response to the prompt.
Correctness Inclusion of all pertinent facts without errors.
Coherence Consistency and clarity of expression.
Complexity Intellectual depth required to write the response (e.g., basic competency vs. domain expertise).
Verbosity Amount of detail included in the response relative to what is asked.
Reasoning–asset generation (corpus construction).
To construct the retrieval corpus, we employ the three reasoning agents ($G_{\text{exp}}$, $G_{\text{diag}}$, $G_{\text{syn}}$) using the o4-mini reasoning model from OpenAI (Achiam et al., 2023). This model is used exclusively for reasoning asset construction rather than downstream inference. All generations follow default API hyperparameters. For each training instance $\tau_{i}$, it produces a reasoning card $\mathcal{C}_{i}$, a diagnostic summary $\mathcal{D}_{i}$, and a set of edit directives $\mathcal{E}_{i}$, which together form the stored reasoning assets $\mathcal{R}_{i}$.
Generation models for evaluation.
For downstream evaluation, we adopt a representative API model GPT-4o (Achiam et al., 2023), and an open-source model LLaMA-3-8B-Instruct (Dubey et al., 2024) as the backbone models. These models are responsible for generating (i) optimized prompts $\hat{p}$ from the Refiner Agent and (ii) the corresponding optimized responses $\hat{r}$ conditioned on $\hat{p}$. Both models are configured with temperature =0=0 to ensure deterministic decoding, and all other hyperparameters remain at their defaults.
Retrieval.
We adopt a sparse retriever based on BM25 (Robertson et al., 2009) over prompt text. For each test prompt $p_{\text{test}}$, we retrieve the top-$k$ training prompts with $k=3$, which we identify as the optimal setting for retrieval [^3] A detailed analysis of $k$ values and their effects on performance is in Section 5.1.1. and attach their paired responses and reasoning assets:
$\{(p_{j},r_{j},\mathcal{R}_{j})\}_{j=1}^{k}.$
4.2. Evaluation Method and Metrics
Judge model.
We employ the ArmoRM-Llama3-8B-v0.1 reward model (Wang et al., 2024b) as an interpretable multi-objective evaluator ($90\%$ benchmark accuracy), which automatically scores response quality. For any prompt-response pair, the judge model returns five HelpSteer-aligned scores (shown in Table 1):
$\mathcal{M}=\{\text{help},\text{corr},\text{coh},\text{comp},\text{verb}\},\quad s_{m}\in[0,4].$
Evaluation Metrics.
Each candidate pair $(p,r)$ is evaluated across the five HelpSteer dimensions $\mathcal{M}=\{\text{help},\text{corr},\text{coh},\text{comp},\text{verb}\}$, where each raw score $s_{m}\in[0,4]$. We normalize scores by dividing by $4$, yielding $\tilde{s}_{m}=s_{m}/4\in[0,1]$. The overall quality score is then defined as the average over all metrics:
$\mathrm{Score}(p,r)=\tfrac{1}{|\mathcal{M}|}\sum_{m\in\mathcal{M}}\tfrac{s_{m}}{4}.$
This produces a single composite value in the range $[0,1]$, reflecting the overall response quality.
4.3. Prompt Design
To guide each agent in MA-SAPO, we carefully designed specialized prompts tailored to their roles in both the training and test phases. [^4] For brevity, we show shortened prompts; full templates are in our code repository.
  1. (1)
    The training prompts (Metric Explainer, Diagnostician, and Action Synthesizer) focus on generating structured reasoning assets by interpreting evaluation scores, diagnosing weaknesses, and synthesizing actionable improvements.
  2. (2)
    The test prompts (Analyzer and Refiner) leverage these assets to contrast retrieved exemplars with the original prompt and iteratively refine it while preserving the user’s intent.
These prompts form the backbone of MA-SAPO’s reasoning-driven optimization pipeline, ensuring interpretability, controllability, and consistency across all stages. The prompt templates are provided below:
Prompt Templates for Different Agents in Training and Testing Phases Metric Explainer Agent Prompt You are an evaluation explainer. Given a user prompt, a model response, and five 0–4 scores (helpfulness, correctness, coherence, complexity, verbosity), write ONE cohesive paragraph (6–9 sentences) that explains WHY each score was assigned and HOW to improve the response. Rules: - Output must be a single paragraph in plain text (no lists, no JSON, no headings) (…) Diagnostician Agent Prompt You are a precise evaluation doctor for LLM outputs. Task: Write ONE cohesive paragraph that (1) diagnoses the main Strength and Weaknesses, (2) explains why they matter in terms of the five metrics, and (3) prescribes concrete fixes. Rules: - Evidence-first: rely only on the prompt + response + reasoning. - Mention each metric once (helpfulness, correctness, coherence, complexity, verbosity) with a brief, specific rationale (…) Output: One paragraph. Action Synthesizer Agent Prompt Mission: Based on the structured diagnosis (Diagnostician Agent), provide a single actionable prompt suggestion that guides improvements while preserving the original intent. Focus on highlighting strengths, addressing weaknesses, and suggesting … Task: Using the diagnosis, propose concise guidance that: 1. Maintains the core topic and original goal of the prompt. 2. Suggests ways to strengthen depth, clarity, and structure. 3. Recommends addition or adjustment (…) Analyzer Agent Prompt You are the Analyzer Agent. Task: (1) Independently evaluate the original prompt (strengths and weaknesses). (2) Analyze retrieved prompt–response–feature triples from three agents. (3) Integrate insights into a unified analysis. Rules: - Begin with direct evaluation of the original prompt. - Evaluate each agent’s output concisely (strengths/weaknesses) (…) Output: A structured but cohesive paragraph-length analysis. Refiner Agent Prompt You are the Refiner Agent. Task: Using Analyzer feedback, optimize the original prompt to resolve all issues while preserving intent. Rules: - Extract the core improvement points from the Analyzer. - Prioritize Action Synthesizer suggestions over others (…) Output: Final optimized prompt.
4.4. Baselines
We implement six baselines under three different categories: single-pass without retrieval (1, 2, and 3), retrieval-augmented without reasoning assets (4), and multi-agent prompt optimization (5 and 6). All methods generate an optimized prompt from the initial prompt, and the final response is produced from that optimized prompt by the same backbone LLM used for the method. Specifically, the six baseline methods are detailed below:
  1. (1)
    Direct Generation: The LLM rewrites the input prompt into an optimized one, without intermediate reasoning.
  2. (2)
    Chain-of-Thought (CoT) [2022] (Wei et al., 2022): The LLM performs step-by-step reasoning about how to improve the prompt, then outputs an optimized version.
  3. (3)
    Role Assignment [2023] (Shanahan et al., 2023): A crafted system role (e.g., “You are a meticulous assistant optimizing prompts for helpfulness, correctness, coherence, complexity, and verbosity.”) guides the LLM to output an optimized prompt in one pass.
  4. (4)
    Retrieval-Augmented Generation (RAG) [2020] (Lewis et al., 2020): M25 retrieves the top-$k=10$ training prompts with their responses and corresponding scores. The LLM references these exemplars to produce an optimized prompt.
  5. (5)
    MAD [2023] (Du et al., 2023): MAD (Multi-Agent Debate) framework organizes several LLM “debaters” that each propose an answer with a rationale, then engage in multi-round critiques of one another’s reasoning. A separate judge (or a consensus rule) evaluates the arguments at each round and selects the final answer, which measurably improves factuality and step-by-step reasoning on strategic-reasoning tasks.
  6. (6)
    MARS [2025] (Zhang et al., 2025): MARS is a hierarchical system of seven agents. A Planner designs the optimization trajectory, while a Teacher, Critic, Student Socratic dialogue iteratively refines the prompt. At each iteration, a Target module evaluates the candidates, logs the history, and the system finally outputs the best prompt.
Table 2. Main results on HelpSteer1 and HelpSteer2. Columns report the five HelpSteer metrics (Help, Corr, Coh, Comp, Verb) normalized to $[0,1]$ and their mean (Avg). Methods are grouped into (i) single-pass prompting, (ii) retrieval-augmented generation (no reasoning assets), and (iii) multi-agent frameworks. Best within each backbone is bold; second best is underlined.
Dataset Methods GPT-4o LLaMA-3-8B
Help Corr Coh Comp Verb Avg Help Corr Coh Comp Verb Avg
HelpSteer1 Direct Generation 0.3216 0.3866 0.7583 0.2951 0.4215 0.4366 0.2549 0.3247 0.7054 0.2702 0.4062 0.3927
Chain-of-Thought (CoT) 0.3223 0.3876 0.7595 0.2935 0.4192 0.4364 0.2359 0.3077 0.6906 0.2627 0.3967 0.3787
Role Assignment 0.3988 0.4679 0.8024 0.3427 0.5008 0.5025 0.2887 0.3516 0.7287 0.3164 0.4517 0.4274
RAG (sparse, k=10k{=}10) 0.3751 0.4402 0.7871 0.3116 0.4779 0.4784 0.2930 0.3594 0.7452 0.3007 0.4377 0.4272
MAD 0.3774 0.4764 0.7954 0.3210 0.5248 0.4990 0.3774 0.5049 0.8067 0.4084 0.6708 0.5531
MARS 0.4159 0.4876 0.7963 0.3293 0.5188 0.5096 0.3901 0.4591 0.7745 0.3512 0.5801 0.5110
MA-SAPO (Ours) 0.5183 0.6260 0.8614 0.5013 0.7363 0.6486 0.4110 0.4868 0.8326 0.4720 0.8433 0.6091
HelpSteer2 Direct Generation 0.3616 0.4700 0.7723 0.3280 0.4913 0.4846 0.2616 0.3636 0.7078 0.2942 0.4421 0.4139
Chain-of-Thought (CoT) 0.2981 0.3958 0.7169 0.2888 0.4709 0.4341 0.1600 0.2545 0.6291 0.2421 0.3812 0.3334
Role Assignment 0.4400 0.5175 0.8221 0.4025 0.5992 0.5563 0.2555 0.3303 0.7050 0.3267 0.4441 0.4123
RAG (sparse, k=10k{=}10) 0.4903 0.5745 0.8642 0.4161 0.6567 0.6003 0.3990 0.4711 0.7989 0.3722 0.5814 0.5245
MAD 0.4167 0.5049 0.8067 0.4084 0.6708 0.5615 0.3971 0.4532 0.7898 0.3998 0.6439 0.5368
MARS 0.4791 0.5569 0.8268 0.4095 0.6234 0.5791 0.4296 0.5019 0.7994 0.3957 0.6181 0.5482
MA-SAPO (Ours) 0.5072 0.6038 0.8527 0.5244 0.7570 0.6490 0.4005 0.4754 0.8294 0.4833 0.8441 0.6065
For all baselines, we adopt GPT-4o and LLaMA-3-8B-Instruct as the backbone models, ensuring consistency and fairness in comparison by aligning with our main experimental setup. Baseline experiments were conducted with the following settings for fairness and fairness of our framework: all LLM temperatures were set to 0 and all other settings were set to LLM defaults. For MAD, it was fixed as $agent=2$ and $round=3$, and MARS designed the planner’s steps in steps 343-4 and set the entire iteration to $5$.
Table 3. Ablation Studies on Retrieval Depth and Test Agent Combination. Varying the number of retrieved training examples ($k$) shows that $k=3$ (our default, denoted as MA-SAPO (k=3)) provides the best balance between contextual coverage and robustness. Collapsing the Analyzer and Refiner into a single agent yields competitive but less consistent results. Best is highlighted in bold; second best is underlined.
Dataset Methods GPT-4o LLaMA-3-8B
Help Corr Coh Comp Verb Avg Help Corr Coh Comp Verb Avg
HelpSteer1 $k=1$ 0.5182 0.6247 0.8543 0.4943 0.7349 0.6453 0.4045 0.4769 0.8268 0.4729 0.8446 0.6051
$k=2$ 0.5211 0.6287 0.8591 0.4960 0.7318 0.6473 0.4002 0.4747 0.8236 0.4726 0.8384 0.6019
$k=4$ 0.5204 0.6266 0.8606 0.4965 0.7369 0.6482 0.4028 0.4778 0.8272 0.4708 0.8389 0.6035
Test Agents Combination 0.4672 0.5415 0.8619 0.3670 0.6403 0.5756 0.4649 0.5203 0.8323 0.4363 0.7249 0.5957
MA-SAPO (k=3) 0.5183 0.6260 0.8614 0.5013 0.7363 0.6486 0.4110 0.4868 0.8326 0.4720 0.8433 0.6091
HelpSteer2 $k=1$ 0.5058 0.5982 0.8458 0.5113 0.7402 0.6403 0.3910 0.4616 0.8215 0.4784 0.8420 0.5989
$k=2$ 0.5107 0.6044 0.8504 0.5158 0.7450 0.6452 0.3980 0.4713 0.8268 0.4845 0.8396 0.6041
$k=4$ 0.5066 0.6012 0.8480 0.5202 0.7451 0.6442 0.3947 0.4685 0.8251 0.4786 0.8395 0.6013
Test Agents Combination 0.5404 0.6012 0.8891 0.4455 0.7398 0.6432 0.4379 0.4720 0.7968 0.4918 0.7321 0.5861
MA-SAPO (k=3) 0.5072 0.6038 0.8527 0.5244 0.7570 0.6490 0.4005 0.4754 0.8294 0.4833 0.8441 0.6065
5. Experimental Results and Analysis
5.1. Main Results
In our main experiments, we compare MA-SAPO with three families of baselines: (1) single-pass prompting methods including Direct Generation, Chain-of-Thought (CoT), and Role Assignment, (2) retrieval-augmented generation (RAG) without reasoning assets, and (3) multi-agent frameworks such as MAD and MARS. As outlined in Section 4.2, we evaluate on both HelpSteer1 and HelpSteer2 using the five annotated quality dimensions and their normalized average score as metrics. The results are shown in Table 2.
In detail, Single-pass methods provide the simplest form of optimization, but their weakness lies in performing edits in isolation. Direct Generation and CoT are essentially one-shot rewrites that lack diagnostic grounding, which makes their improvements brittle and often inconsistent across metrics. Role Assignment can guide the model to consider multiple aspects explicitly, yet it still suffers from over-reliance on handcrafted role descriptions and cannot adapt when those roles fail to cover task-specific subtleties. RAG improves upon these by injecting exemplars, but the optimization process remains unguided: retrieved examples are consumed without explanation of why they are useful, leaving the refinement process heuristic and prone to inconsistency. As a result, these methods either underperform on complex prompts or achieve higher scores only at the cost of verbosity and drift in task semantics.
Recent multi-agent frameworks such as MAD and MARS perform better by coordinating multiple roles in a structured workflow, yet they remain outcome-driven and heavy in cost [^5] A detailed cost and latency comparisons are provided in Section 5.4.. Their reliance on repeated debate or critique cycles incurs significant computational overhead, while still treating evaluation as a scalar target rather than a source of actionable reasoning. In contrast, our MA-SAPO consistently surpasses all baselines across datasets and backbones by explicitly converting evaluation signals into reusable reasoning assets. The Analyzer to Refiner pipeline applies only evidence-backed edits drawn from these assets, ensuring that optimizations are not only effective but also interpretable and controllable [^6] A detailed case study is provided in Section 5.5.. For instance, on HelpSteer1 with GPT-4o, MA-SAPO achieves 0.6486 average score compared to 0.5096 for MARS and 0.4784 for RAG. Similar margins hold across HelpSteer2 and with LLaMA-3-8B, confirming that structured reasoning, rather than trial-and-error search, is the main driver of performance. Overall, MA-SAPO combines accuracy and efficiency, making it a more practical and principled framework for prompt optimization than competitive baselines.
5.2. Ablation Study
5.2.1. Top-k Retrieval Variations
To examine how the number of retrieved training examples influences the performance of MA-SAPO, we vary the retrieval depth $k\in\{1,2,3,4\}$. Table 3 reports results for both GPT-4o and LLaMA-3-8B models across HelpSteer1 and HelpSteer2.
For both backbones, increasing $k$ from 1 to 3 generally improves overall performance. With $k=1$, the Analyzer is under-informed, as it can only compare the test prompt against a single exemplar, limiting the diversity of reasoning assets available. Moving to $k=2$ provides broader evidence and modest gains, while $k=3$ yields the best balance across all five metrics and achieves the highest average performance. At $k=4$, results remain competitive but show slight declines in some dimensions, suggesting that introducing too many exemplars adds noise and creates conflicting diagnostic signals.
These findings highlight a clear trade-off: smaller $k$ restricts contextual diversity, while larger $k$ risks diluting the relevance of retrieved assets. Overall, $k=3$ achieves the most consistent and robust performance across datasets and backbones, and we adopt it as the default retrieval setting in all subsequent experiments.
5.2.2. Test Agents Combination
To further assess the robustness of MA-SAPO, we conduct an ablation study where the two test-phase agents are merged into a single combined agent, referred to as Test Agents Combination. Table 3 summarizes the results for both GPT-4o and LLaMA-3-8B on HelpSteer1 and HelpSteer2.
We observe that the combined variant achieves competitive performance, occasionally surpassing the full MA-SAPO in a single metric. However, the average performance consistently falls behind the full framework. This suggests that while a unified agent can capture certain improvements, it lacks the structured diagnostic pipeline of MA-SAPO, where the Analyzer first produces evidence-grounded reports and the Refiner then executes targeted edits.
These findings demonstrate the necessity of separating the two roles in the test phase. The Analyzer to Refiner pipeline enforces a more interpretable and controllable workflow, ensuring that refinements are justified by retrieved reasoning assets rather than introduced implicitly by a single agent. As a result, the modular design of MA-SAPO is not only more reliable but also essential for achieving consistent gains across datasets and backbones.
5.3. Qualitative Analysis
To complement the automatic benchmark results, we conducted a human evaluation designed to assess both the interpretability and controllability of MA-SAPO’s optimization process. This evaluation focused on two complementary hypotheses: (H1) that multi-agent reasoning produces higher-quality reasoning than a single-agent baseline, and (H2) that MA-SAPO preserves the semantic direction and intent of the original prompts during optimization.
All evaluations were conducted by 14 independent human annotators (mean age = 29.1, SD = 3.4), who were either data scientists, AI engineers, or graduate students majoring in data science or related fields. Each annotator was familiar with language model evaluation and prompt optimization procedures. The evaluation employed a 5-point Likert scale for reasoning quality (H1) and a 4-point scale for directional consistency (H2). The following subsections describe the evaluation setup and results.
5.3.1. Evaluating Multi-Agent Reasoning Quality (H1)
To assess whether MA-SAPO enhances reasoning quality, we randomly sampled 30 reasoning cases from the HelpSteer1 training set. Each case included two reasoning outputs: one generated by a single-agent baseline and one by MA-SAPO’s multi-agent reasoning pipeline, resulting in 60 reasoning outputs in total. These outputs were independently evaluated by 14 human annotators, yielding 840 individual evaluations (30 cases $\times$ 2 outputs $\times$ 14 annotators). Each output was rated along three qualitative dimensions:
  • Usefulness: How useful the response is in addressing the prompt.
  • Accuracy: How factually correct the reasoning content is.
  • Consistency: How coherent and logically structured the writing is.
Each dimension was rated on a scale from 1 (very low) to 5 (very high).
Results.
As summarized in Table 4, MA-SAPO outperformed the single-agent baseline across all three dimensions. The largest gains were observed in perceived usefulness and factual accuracy, both showing statistically significant improvements (*p* ¡ 0.05, paired $t$-test, $n=30$). These findings indicate that multi-agent reasoning yields more helpful, accurate, and consistent responses.
Table 4. Human evaluation of reasoning quality. MA-SAPO outperforms the single-agent baseline in usefulness, factual accuracy, and consistency. Scores are averaged on a 1–5 scale; asterisks denote significance levels from paired $t$-tests ($^{*}p<0.05$, $n=30$).
Method Usefulness Accuracy Consistency Mean
Single-Agent 3.64 3.63 3.81 3.69
MA-SAPO (Ours) 3.89 3.87 4.02 3.93
$\Delta$ (Improvement) +0.25 +0.24 +0.21 +0.23
5.3.2. Evaluating Directional Consistency of Optimized Prompts (H2)
To evaluate whether MA-SAPO preserves the semantic intent of prompts during optimization, we randomly sampled 40 prompt pairs, each consisting of an original prompt ($A$) and its optimized version ($B$) generated by MA-SAPO. Each pair was independently rated by 14 human annotators, resulting in a total of 560 evaluations (40 pairs $\times$ 14 annotators). Annotators rated the directional consistency of each pair on a 4-point Likert scale (1 = completely changed, 2 = partially changed, 3 = mostly preserved, 4 = fully preserved), indicating how well the optimized prompt maintained the intent, goal, and domain of the original.
Results.
Figure 3 visualizes the distribution of annotator ratings for directional consistency. Most scores are concentrated between 3 and 4, indicating that the optimized prompts largely preserved the original semantic intent. The mean rating was 3.36 with a standard deviation of 0.74 ($n=560$), demonstrating that MA-SAPO effectively enhances prompt clarity and structure while maintaining semantic stability. The relatively moderate standard deviation further suggests that annotators consistently judged the optimized prompts as semantically aligned with their originals.
Figure 3. Distribution of directional consistency ratings (n=560). Ratings on a 4-point scale (1 = changed, 4 = preserved). The dashed line marks the mean (3.36) and the shaded band shows one standard deviation (SD = 0.74).
5.3.3. Summary of Human Evaluation
The human evaluation results highlight two complementary aspects of MA-SAPO. First, the multi-agent reasoning framework generates responses that are more useful, accurate, and consistent than those from a single-agent baseline, confirming the advantage of role-specialized reasoning. Second, MA-SAPO maintains strong semantic stability during prompt optimization, ensuring that improved clarity and structure do not distort the original task intent. Together, these findings demonstrate that MA-SAPO achieves both reasoning quality and semantic controllability, reinforcing the quantitative evidence presented in Section 5.3.
5.4. Cost and Latency
Table 5. Cost and Latency Comparison. MA-SAPO requires far fewer tokens and calls than other competitive multi-agent baselines, while maintaining superior optimization performance. Lower is better ($\downarrow$).
Method # Input Tokens $\downarrow$ # Output Tokens $\downarrow$ # Calls $\downarrow$
MAD 5,217 2,170 8
MARS 17,149 6,789 60
MA-SAPO (Ours) 4,968 526 2
To complement the effectiveness results, we also compare the computational cost of different multi-agent frameworks. Table 5 reports the average number of input tokens, output tokens, and calls per instance. We observe that existing multi-agent systems such as MAD and MARS incur substantial overhead. MAD requires around 5K input tokens, 2K output tokens, and 8 calls on average due to repeated multi-round debates. MARS is even heavier, consuming more than 17K input tokens, nearly 7K output tokens, and about 60 calls per instance, reflecting its iterative planner–critic–student workflow. In contrast, MA-SAPO is significantly more efficient: by constructing reasoning assets offline and adopting a single Analyzer to Refiner loop at test time, it reduces the runtime cost to roughly 5K input tokens, 0.5K output tokens, and only 2 calls per instance.
These results highlight that MA-SAPO not only improves optimization quality but also achieves a more favorable cost–latency trade-off. While MAD and MARS achieve moderate gains at the expense of heavy resource usage, MA-SAPO provides consistent improvements while keeping computational demands.
Figure 4. Case study illustrating MA-SAPO’s multi-agent prompt optimization process. Starting from an initial user prompt and evaluation scores, the framework retrieves reasoning assets, analyzes weaknesses, and iteratively refines the prompt. The regenerated response shows improved clarity, structure, and contextual depth compared to the original.
5.5. Case Study
To further illustrate the effectiveness and practical utility of MA-SAPO, we present a case study drawn from the HelpSteer2 validation dataset. [^7] Additional case studies demonstrating diverse application scenarios are available in our code repository. This example highlights MA-SAPO’s strengths in collaborative reasoning, context-aware prompt refinement, and robust adaptation across diverse input types, thereby providing concrete evidence of the framework’s ability to enhance prompt optimization in complex reasoning scenarios.
The case shown in Figure 4 demonstrates the end-to-end prompt optimization process. It begins with an initial input prompt, “What are 10 primary ways ChatGPT will change 4th–12th grade education in the United States?”, alongside the corresponding model response and evaluation scores. MA-SAPO then dynamically retrieves three semantically and thematically relevant exemplars from the training corpus (e.g., “applying ChatGPT in language learning and teaching”). These retrieved assets introduce diverse educational contexts and prompt formulations, implicitly guiding refinements such as narrowing task scope and aligning conceptual goals with actionable applications.
The Analyzer Agent identifies weaknesses in the original prompt notably, vague task specification and overly broad phrasing-both of which risk eliciting generic responses. Building on this analysis, the Refiner Agent reformulates the prompt to preserve its educational intent while incorporating a more actionable subtask, such as exploring business opportunities tied to ChatGPT-based assessment practices. The optimized prompt exhibits clearer structure, explicit segmentation, and richer contextual depth. When regenerated with the refined prompt, the model produces a response that is more coherent, pedagogically grounded, and practically insightful.
Overall, this case underscores how MA-SAPO integrates multi-agent reasoning and retrieved exemplars to achieve superior prompt optimization, outperforming conventional single-agent or static fine-tuning approaches in terms of coherence, interpretability, and alignment with user intent.
6. Conclusion
We propose MA-SAPO, a novel multi-agent framework for interpretable and controllable prompt optimization. Compared to prior methods that either perform one-shot rewrites, rely on unguided retrieval, or employ heavy multi-agent debates, MA-SAPO explicitly transforms evaluation outcomes into reusable reasoning assets and grounds refinements in evidence-backed Analyzer to Refiner interactions. Our contributions extend beyond framework design to include comprehensive experiments on two benchmarks and multiple backbones, ablation studies on retrieval depth and agent configuration, and human evaluations of reasoning quality and semantic consistency. Extensive results demonstrate that MA-SAPO consistently outperforms existing approaches, achieving higher optimization quality while remaining efficient in cost and latency.
Despite these advantages, MA-SAPO currently depends on the quality of human-annotated scores and a fixed semi-structured schema for asset construction, while also relying on a sparse BM25 retriever and a single reward model-factors that may overlook semantically relevant exemplars and introduce evaluator bias. In future work, we plan to improve robustness by integrating a dedicated Feedback Agent to validate asset quality and mitigate bias, as well as by developing hybrid dense–sparse retrieval strategies. We also aim to extend the framework to multi-turn prompts to enable more comprehensive evaluation.

### v2

@@
-Prompt Optimization via Retrieved Reasoning Assets and Multi-Agent Analysis
-Abstract.
-Prompt optimization has emerged as an effective alternative to retraining for improving the performance of Large Language Models (LLMs).
-However, most existing approaches treat evaluation as a black box, relying solely on numerical scores while offering limited insight into why a prompt succeeds or fails.
-They also depend heavily on trial-and-error refinements, which are difficult to interpret and control.
-In this paper, we introduce MA-SAPO, a Multi-Agent framework for Score-Aware Prompt Optimization.
-Compared to prior methods, MA-SAPO explicitly couples evaluation outcomes with structured reasoning to guide systematic edits.
-The framework specifically consists of two stages: during the Reasoning Phase, agents collaboratively explain metric scores, diagnose weaknesses, and synthesize targeted refinements that are stored as reusable reasoning assets; during the Test Phase, agents retrieve these assets to analyze optimized prompts and apply only evidence-grounded edits.
-By turning evaluation signals into interpretable reasoning chains, MA-SAPO produces prompt refinements that are more transparent, auditable, and controllable.
-Experiments on the HelpSteer1/2 benchmarks demonstrate consistent improvements over single-pass prompting, retrieval-augmented baselines, and prior multi-agent strategies, validating the effectiveness of our approach.
-[^1] Code is available at: https://anonymous.4open.science/r/MA-SAPO-F313.
-Automatic Prompt Optimization, Score-Aware Prompt Optimization, Retrieval-Augmented Generation, Large Language Models (LLMs), Multi-Agent System
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Multi-agent systemsccs: Information systems Query representationccs: Computing methodologies Natural language generationccs: Information systems Relevance assessment
-Figure 1. Comparative overview of prompt optimization methods. MA-SAPO enhances interpretability by linking evaluation scores to reasoning-driven and evidence-based refinements.
-1. Introduction
+MA-SAPO: Multi-Agent Reasoning for Score-Aware Prompt Optimization
+Abstract
+Prompt optimization has become a practical way to improve the performance of Large Language Models (LLMs) without retraining.
+However, most existing frameworks treat evaluation as a black box, relying solely on outcome scores without explaining why prompts succeed or fail.
+Moreover, they involve repetitive trial-and-error refinements that remain implicit, offering limited interpretability or actionable guidance for systematic improvement.
+In this paper, we propose MA-SAPO: a new Multi-Agent Reasoning for Score Aware Prompt Optimization framework that links evaluation outcomes directly to targeted refinements.
+Specifically, in the Training Phase, multiple agents interpret evaluation scores, diagnose weaknesses, and generate concrete revision directives, which are stored as reusable reasoning assets.
+In the Test Phase, an analyzer agent retrieves relevant exemplars and assets for a new prompt, and a refiner agent applies evidence-based edits to improve the prompt and its response.
+By grounding optimization in structured reasoning, MA-SAPO ensures edits are interpretable, auditable, and controllable.
+Experiments on the HelpSteer1/2 benchmarks show that our framework consistently outperforms single-pass prompting, retrieval-augmented generation, and prior multi-agent methods across multiple evaluation metrics.
+1 Introduction
 Large Language Models (LLMs) have emerged as powerful tools capable of tackling a wide range of tasks, from reasoning and summarization to complex dialogue and code generation (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023).
-However, their performance remains highly sensitive to the wording, structure, and examples embedded within prompts (Pryzant et al., 2023; He et al., 2024; Errica et al., 2024; Cao et al., 2024; Razavi et al., 2025).
-Consequently, prompt optimization has rapidly emerged as a practical alternative for improving model behavior, as it avoids the need for costly retraining or parameter updates.
-Early research explored single-pass prompt optimization strategies that treated the model itself as the optimizer.
-For instance, methods such as Chain-of-Thought (CoT) prompting (Wei et al., 2022), role assignment through carefully designed system instructions (Wu et al., 2023; Kong et al., 2023), or structured variants such as Tree-of-Thought (ToT) (Yao et al., 2023), Graph-of-Thought (GoT) (Besta et al., 2024), and step-back prompting (Zheng et al., 2023) demonstrated that reasoning patterns or role signals embedded into prompts could substantially improve downstream performance.
-While effective, these approaches typically operate in isolation: a single model instance refines prompts without incorporating systematic feedback from prior attempts or external resources.
-As LLMs have advanced, multi-agent frameworks have gained traction as a way to diversify perspectives and coordinate multiple roles in the optimization process (Zhou et al., 2025; Zhang et al., 2025; Han et al., 2025).
-By assigning diverse specialized agents to different tasks such as decomposition, evaluation, or refinement, these systems move beyond single-pass rewriting to more structured optimization pipelines.
-At the same time, optimization-theoretic methods (Li et al., 2025b; Cui et al., 2025) and label-efficient pipelines (Xiang et al., 2025; Dong et al., 2025) have sought to reduce the reliance on costly feedback while ensuring scalability across tasks.
-Despite these advances, several common limitations remain: (1) most existing frameworks reduce evaluation to outcome scores, treating it as a black box and leaving practitioners uncertain about why certain prompts succeed or fail; (2) optimization often relies on repetitive trial-and-error refinements that consume significant computation while offering little transparency; (3) even when reasoning is incorporated, it usually remains implicit and is not transformed into explicit, auditable artifacts that can guide systematic edits; (4) current pipelines largely optimize for higher outcome scores but offer limited interpretability and controllability, preventing users from understanding trade-offs or applying targeted adjustments.
-To address these limitations, we propose MA-SAPO, a sequential reasoning framework for prompt optimization that explicitly links metric outcomes to actionable edits through structured reasoning artifacts.
-Specifically, MA-SAPO operates in two complementary phases.
-In the reasoning phase, three agents transform annotated scores into progressively enriched artifacts: (1) a Metric Explainer Agent interprets evaluation dimensions, (2) a Diagnostician Agent uncovers error sources and trade-offs, and (3) an Action Synthesizer Agent produces concrete edit directives.
-These artifacts are stored as reusable reasoning assets.
-In the test phase, a retrieval-augmented pipeline leverages these assets: (1) an Analyzer Agent contrasts the current prompt with retrieved exemplars to identify improvement opportunities, while (2) a Refiner Agent applies targeted edits supported by diagnostic evidence, ensuring that optimizations are interpretable and controllable.
-Extensive experiments on the open-source HelpSteer1/2 benchmarks (Wang et al., 2023, 2024a) demonstrate that MA-SAPO consistently outperforms single-pass prompting, retrieval-augmented generation, and prior multi-agent baselines, highlighting the importance of structured reasoning as a bridge between evaluation and optimization.
-Our main contributions include:
-  • We propose a score-aware multi-agent training pipeline that distills evaluation outcomes into reusable, semi-structured reasoning assets, enabling transparent prompt explanations, fine-grained diagnoses, and actionable edit directives.
-  • We design MA-SAPO, a modular and computationally efficient framework that achieves interpretable, auditable, and controllable improvements across diverse datasets and model backbones, while substantially reducing token and API call budgets.
-  • We conduct extensive experiments on HelpSteer1 and HelpSteer2 benchmarks, demonstrating consistent gains across multiple evaluation metrics and outperforming six state-of-the-art baselines in both effectiveness and efficiency.
-2. Related Work
-Single-Pass Prompt Optimization.
-Prompt optimization has rapidly evolved as a practical means of improving the performance of LLMs without retraining (Wei et al., 2022; Pryzant et al., 2023; Sun et al., 2023; Li et al., 2025a).
-Early studies highlighted the centrality of prompts as control variables for steering model behavior.
-Specifically, few-shot prompting (Brown et al., 2020) demonstrated that models could adapt to unseen tasks using in-context examples alone, while retrieval-augmented generation (RAG) (Lewis et al., 2020; Lee et al., 2025) injected external knowledge to improve factuality and adaptability.
-These two paradigms, context construction and evidence injection, laid the foundation for single-pass prompt optimization methods.
-Subsequent research explored reasoning-augmented prompting.
-Chain-of-Thought (CoT) (Wei et al., 2022) introduced intermediate reasoning steps.
-Building on this idea, Tree-of-Thought (ToT) (Yao et al., 2023) generalized linear reasoning chains into branching structures with self-evaluation via lookahead and backtracking.
-Other structured strategies, such as Skeleton-of-Thought (SoT) (Ning et al., 2023), separate planning from execution to parallelize generation, while Graph-of-Thought (GoT) (Besta et al., 2024) organizes reasoning trajectories into graph topologies.
-Evolutionary and heuristic single-pass approaches have also been explored.
-For instance, PromptBreeder (Fernando et al., 2023) evolves a population of prompts through mutation and selection, while PromptWizard (Agarwal et al., 2024) refines prompts via feedback-driven critique-and-synthesis loops.
-Multi-Agent Based Prompt Optimization.
-Beyond single-pass methods, multi-agent systems have emerged as a powerful paradigm for prompt optimization (Du et al., 2023; Liu et al., 2024; Chen et al., 2024; Li et al., 2024).
-Early agentic frameworks such as Self-Ask (Press et al., 2022) introduced decomposition by generating sub-questions and integrating external retrieval, laying the groundwork for cooperative prompting strategies.
-Also, Multi-Agent Debate (MAD) (Du et al., 2023) advanced this idea by allowing agents to iteratively present and critique answers in a roundtable format, thereby improving factual grounding and reasoning quality.
-Reflection-based frameworks (Shinn et al., 2023; Khan et al., 2024) further enhanced adaptability by storing feedback as memory for subsequent optimization attempts.
-Building on this foundation, recent systems formalized agent roles and workflows to optimize prompts more systematically (Li et al., 2023).
-Subsequent work explicitly targeted prompt optimization as the objective.
-For instance, MASS (Zhou et al., 2025) searches over agent topologies while performing block-level prompt refinement, followed by global optimization.
-Additionally, MARS (Zhang et al., 2025) leverages a Planner and Teacher–Critic–Student dialogue to iteratively propose, critique, and revise prompts in structured rounds, and also MAPGD (Han et al., 2025) integrates pseudo-gradient feedback with bandit-style exploration, offering sample-efficient optimization through agent collaboration.
-However, most existing frameworks treat evaluation as a black box with limited interpretability and rely on costly trial-and-error refinements.
-MA-SAPO addresses this by grounding optimization in score-aware reasoning assets that link evaluation outcomes to actionable, interpretable, and auditable refinements.
-Figure 2.
-Overview of the MA-SAPO framework.
-The training phase constructs reasoning assets $(\mathcal{R}_{i}=(\mathcal{C}_{i},\mathcal{D}_{i},\mathcal{E}_{i}))$ from annotated prompt–response pairs, while the test phase retrieves top-k training examples and their reasoning assets to refine a new test prompt.
-3. Methodology
-3.1. Preliminaries
-We denote the training dataset as
+However, their performance remains highly sensitive to the wording and examples embedded within prompts (Pryzant et al., 2023; He et al., 2024; Cao et al., 2024; Razavi et al., 2025).
+Consequently, prompt optimization has rapidly emerged as a practical alternative to costly retraining or parameter updates for improving model behavior.
+Early research explored single-pass prompt optimization that effectively treats the LLM as its own optimizer.
+Approaches such as Chain-of-Thought (CoT) prompting (Wei et al., 2022), role conditioning via carefully crafted system instructions (Wu et al., 2023; Kong et al., 2024), and structured reasoning schemes including Tree-of-Thought (ToT) (Yao et al., 2023), Graph-of-Thought (GoT) (Besta et al., 2024), and step-back prompting (Zheng et al., 2024) demonstrated that encoding explicit reasoning structure or role signals in the prompt can materially improve downstream performance.
+Nevertheless, these methods are largely self-contained, relying on a single model instance to revise prompts without leveraging systematic feedback across iterations or external resources.
+As LLMs have advanced, multi-agent frameworks have emerged to broaden viewpoints and coordinate specialized roles in prompt optimization (Zhou et al., 2025; Zhang et al., 2025; Han et al., 2025).
+By delegating subtasks such as decomposition, evaluation, and refinement to distinct agents, these systems extend beyond single-pass rewriting into more structured pipelines.
+In parallel, optimization-theoretic approaches (Li et al., 2025a; Cui et al., 2025) and label-efficient schemes (Xiang et al., 2025; Dong et al., 2025) aim to curb reliance on expensive feedback while scaling across tasks.
+Despite these advances, several common limitations remain: (i) evaluation is often collapsed to scalar scores, obscuring why prompts succeed or fail; (ii) refinement remains largely trial-and-error, incurring high computational cost with limited transparency; (iii) reasoning, when used, is typically implicit rather than distilled into auditable artifacts that drive targeted edits; and (iv) pipelines prioritize score gains over interpretability and controllability, which hinders practitioners from analyzing trade-offs and making precise adjustments in real-world applications.
+To address these limitations, we introduce MA-SAPO, a sequential reasoning framework for prompt optimization that explicitly maps metric outcomes to actionable edits via structured reasoning artifacts.
+MA-SAPO has two phases.
+In the training phase, three agents convert annotated scores into reusable assets: (1) a Metric Explainer interprets evaluation dimensions, (2) a Diagnostician identifies error sources and trade-offs, and (3) an Action Synthesizer produces concrete edit directives.
+In the test phase, a retrieval-augmented pipeline applies these assets: an Analyzer contrasts the current prompt with retrieved exemplars to surface improvement opportunities, and a Refiner performs targeted edits grounded in diagnostic evidence, yielding interpretable and controllable optimization.
+Experiments on HelpSteer1/2 (Wang et al., 2024c; b) show that MA-SAPO consistently surpasses single-pass prompting, retrieval-augmented generation, and prior multi-agent baselines, underscoring structured reasoning as an effective bridge between evaluation and optimization.
+Our contributions are:
+  • A score-aware training pipeline that distills evaluation outcomes into reusable, semi-structured reasoning assets for explanation, diagnosis, and edit directives.
+  • MA-SAPO, a modular and compute-efficient framework that delivers interpretable, auditable, and controllable gains while reducing token and API-calls.
+  • Extensive results on HelpSteer1/2 demonstrating consistent multi-metric improvements and outperforming six baselines in both effectiveness and efficiency.
+2 Related work
+2.1 Single-pass prompt optimization
+Prompt optimization has become an effective way to improve LLM performance without retraining (Wei et al., 2022; Pryzant et al., 2023; Sun et al., 2023; Li et al., 2025b).
+Early work established prompts as key control variables for steering model behavior.
+In particular, few-shot prompting (Brown et al., 2020) showed that models can adapt to new tasks through in-context examples, while retrieval-augmented generation (RAG) (Lewis et al., 2020; Lee et al., 2025) improves factuality by injecting external knowledge.
+Together, these approaches laid the foundation for single-pass prompt optimization.
+Later work introduced reasoning-augmented prompting.
+Chain-of-Thought (CoT) (Wei et al., 2022) added intermediate reasoning steps, and Tree-of-Thought (ToT) (Yao et al., 2023) extended this into branching reasoning with self-evaluation.
+Other structured variants include Skeleton-of-Thought (SoT) (Ning et al., 2023), which separates planning from execution, and Graph-of-Thought (GoT) (Besta et al., 2024), which models reasoning as graph-structured trajectories.
+Evolutionary and heuristic methods have also been explored: PromptBreeder (Fernando et al., 2023) evolves prompts through mutation and selection, while PromptWizard (Agarwal et al., 2024) refines them via critique-and-synthesis loops.
+2.2 Multi-agent based prompt optimization
+Beyond single-pass methods, multi-agent systems have become an important paradigm for prompt optimization (Du et al., 2024; Liu et al., 2024; Chen et al., 2024; Li et al., 2024; Seo et al., 2026).
+Early agentic approaches such as Self-Ask (Press et al., 2023) introduced question decomposition and retrieval, while Multi-Agent Debate (MAD) (Du et al., 2024) improved reasoning through iterative answer exchange and critique.
+Reflection-based methods (Shinn et al., 2023; Khan et al., 2024) further enhanced adaptability by incorporating feedback from prior attempts.
+Building on these ideas, later frameworks structured agent roles and workflows more explicitly (Li et al., 2023) and began targeting prompt optimization itself.
+For example, MASS (Zhou et al., 2025) searches over agent topologies with local and global prompt refinement, MARS (Zhang et al., 2025) uses Planner and Teacher–Critic–Student interactions for iterative revision, and MAPGD (Han et al., 2025) combines pseudo-gradient feedback with bandit-based exploration.
+However, many existing methods still treat evaluation as a black box and rely on costly trial-and-error refinement.
+MA-SAPO instead grounds optimization in score-aware reasoning assets, making revisions more interpretable, auditable, and actionable.
+3 Methodology
+Figure 1: MA-SAPO pipeline overview.
+Three training agents convert annotated prompt–response pairs into reusable Reasoning Assets $\mathcal{R}_{i}=(\mathcal{C}_{i},\mathcal{D}_{i},\mathcal{E}_{i})$.
+At test time, retrieved assets guide the Analyzer $G_{\mathrm{ana}}$ and Refiner $G_{\mathrm{ref}}$ to produce an optimized prompt $\hat{p}$ with higher scores across all five evaluation dimensions.
+3.1 Preliminaries
+Let the training dataset be denoted as
 $\mathcal{D}_{train}=\{\tau_{i}\}_{i=1}^{N},\quad\tau_{i}=(p_{i},r_{i},\mathcal{S}_{i}),$
@@
 where only prompts are available, and corresponding responses are generated during evaluation.
-Our objective is to generate structured reasoning assets from $\mathcal{D}_{train}$ that can later be retrieved and utilized to optimize new prompts and their responses.
+Our objective is to generate structured reasoning assets from $\mathcal{D}_{train}$ that can be retrieved to optimize new prompts and their corresponding responses.
 Thus, the training phase serves as data generation, while the test phase operates as a retrieval-augmented optimization pipeline.
@@
 The overview of the MA-SAPO framework is shown in Figure 1.
-3.2. Training Phase: Reasoning Asset Construction
+3.2 Training phase: reasoning asset construction
 In the training phase, three agents are executed sequentially for each $\tau_{i}\in\mathcal{D}_{train}$.
 All three agents operate on the same triplet $\tau_{i}=(p_{i},r_{i},\mathcal{S}_{i})$, ensuring that their reasoning is consistently grounded in both the input and output of the original instance.
-3.2.1. Metric Explainer Agent $G_{\text{exp}}$
-The Metric Explainer Agent provides natural language justifications for the given scores. Formally, given $\tau_{i}$, the agent produces a reasoning card $\mathcal{C}_{i}$:
-(1) $\mathcal{C}_{i}=G_{\text{exp}}(\tau_{i},P_{\text{exp}}),$
+3.2.1 Metric explainer agent $G_{\text{exp}}$
+The Metric Explainer Agent generates natural language justifications for the assigned scores. Formally, given $\tau_{i}$, the agent produces a reasoning card $\mathcal{C}_{i}$:
+$\mathcal{C}_{i}=G_{\text{exp}}(\tau_{i},P_{\text{exp}}),$ (1)
 where $P_{\text{exp}}$ is the system prompt guiding explanation.
@@
 The reasoning card $\mathcal{C}_{i}$ then serves as input to the next agent, together with $\tau_{i}$.
-3.2.2. Diagnostician Agent $G_{\text{diag}}$
+3.2.2 Diagnostician agent $G_{\text{diag}}$
 The Diagnostician Agent operates on $\tau_{i}$ and the reasoning card $\mathcal{C}_{i}$.
@@
 coherence).
-(2) $\mathcal{D}_{i}=G_{\text{diag}}(\tau_{i},\mathcal{C}_{i},P_{\text{diag}}).$
+$\mathcal{D}_{i}=G_{\text{diag}}(\tau_{i},\mathcal{C}_{i},P_{\text{diag}}).$ (2)
 The diagnostic summary $\mathcal{D}_{i}$ is then passed to the next stage, along with $\tau_{i}$.
-3.2.3. Action Synthesizer Agent $G_{\text{syn}}$
+3.2.3 Action synthesizer agent $G_{\text{syn}}$
 The Action Synthesizer Agent receives $\tau_{i}$ together with $\mathcal{C}_{i}$ and $\mathcal{D}_{i}$.
@@
 Each directive corresponds to a concrete modification strategy for improving $\tau_{i}$:
-(3) $\mathcal{E}_{i}=G_{\text{syn}}(\tau_{i},\mathcal{C}_{i},\mathcal{D}_{i},P_{\text{syn}}),$
+$\mathcal{E}_{i}=G_{\text{syn}}(\tau_{i},\mathcal{C}_{i},\mathcal{D}_{i},P_{\text{syn}}),$ (3)
 where $\mathcal{E}_{i}=\{e_{1},e_{2},\dots,e_{m}\}$ is a set of recommended edits.
@@
 $\mathcal{R}_{i}=(\mathcal{C}_{i},\mathcal{D}_{i},\mathcal{E}_{i}).$
-As a result, $\mathcal{R}_{i}$ is stored as the reasoning assets aligned with $\tau_{i}$, forming the retrieval corpus for the test phase.
+As a result, $\mathcal{R}_{i}$ is stored as reasoning assets aligned with $\tau_{i}$, forming the retrieval corpus for the test phase.
 Importantly, these reasoning assets are designed as semi-structured text, which makes them machine-parseable and directly usable by downstream agents in the test phase.
-3.3. Test Phase: Retrieval and Reformulation
+For instance, edit directives often follow a pattern such as “add a constraint to …”, which the refiner agent can reliably identify and execute.
+3.3 Test phase: retrieval and reformulation
 In the test phase, optimization is performed via retrieval-augmented generation.
@@
 $\{(\tau_{j},\mathcal{R}_{j})\}_{j=1}^{k}.$
-3.3.1. Analyzer Agent $G_{\text{ana}}$
+3.3.1 Analyzer agent $G_{\text{ana}}$
 The Analyzer Agent compares $p_{\text{test}}$ against retrieved examples to identify improvement points. It outputs an improvement report $\mathcal{A}$:
-(4) $\mathcal{A}=G_{\text{ana}}(p_{\text{test}},\{\tau_{i},\mathcal{R}_{j}\}_{j=1}^{k},P_{ana}).$
+$\mathcal{A}=G_{\text{ana}}(p_{\text{test}},\{\tau_{i},\mathcal{R}_{j}\}_{j=1}^{k},P_{\text{ana}}).$ (4)
 The Analyzer Agent transforms raw retrieval into structured insights rather than merely transferring edits from similar prompts.
 The agent highlights concrete weaknesses and improvement opportunities by contrasting the test prompt with retrieved reasoning assets.
-3.3.2. Refiner Agent $G_{\text{ref}}$
+3.3.2 Refiner agent $G_{\text{ref}}$
 The Refiner Agent regenerates an optimized prompt $\hat{p}$ by incorporating the improvement report:
-(5) $\hat{p}=G_{\text{ref}}(p_{\text{test}},\mathcal{A},P_{\text{ref}}).$
+$\hat{p}=G_{\text{ref}}(p_{\text{test}},\mathcal{A},P_{\text{ref}}).$ (5)
 It operationalizes the analyzer’s insights into a concrete optimized prompt.
@@
 The final optimized prompt $\hat{p}$ is then used to generate a corresponding optimized response $\hat{r}$.
-3.4. Evaluation
-[…]
-(6) $\text{Score}(\hat{p},\hat{r})=\langle s_{help},s_{corr},s_{coh},s_{comp},s_{verb}\rangle,$
-where each $s_{m}\in[0,4]$ for $m\in\mathcal{M}$.
-These scores are normalized into the $[0,1]$ range and then aggregated into a single composite value.
-[^2] This evaluation methodology is implemented in our setup in Section 4.2.
-Additionally, We complement automatic metrics with a human evaluation (Section 5.3) that tests (H1) the usefulness, accuracy, and consistency of the reasoning and (H2) whether optimized prompts preserve the original intent; 14 expert annotators, all data scientists or AI engineers, participated.
-4. Experimental Design and Setup
-4.1. Datasets
-We conduct our experiments on the HelpSteer family of datasets (Wang et al., 2023, 2024a), which provide human-annotated prompt-response pairs for evaluating prompt optimization in Large Language Models (LLMs).
-Both datasets share the same annotation schema across five quality dimensions, but differ in size and design.
-We specifically build the retrieval corpus from the HelpSteer2 training dataset, where each instance $\tau_{i}=(p_{i},r_{i},\mathcal{S}_{i})$ is augmented with the reasoning assets $\mathcal{R}_{i}=(\mathcal{C}_{i},\mathcal{D}_{i},\mathcal{E}_{i})$ produced during our training pipeline.
-Evaluation is conducted on both HelpSteer1 and HelpSteer2 validation prompts.
-HelpSteer1 contains 35.3k training and .79k validation samples, each comprising a prompt, a response, and five human-annotated attributes scored on a 0-4 scale.
-It served as the first large-scale benchmark for aligning model outputs with human feedback and provides broad coverage of prompt-response styles.
-HelpSteer2 is a newer dataset containing 20.3k training and 1.04k validation samples, with higher-quality annotations, multi-turn prompts (about 29% of cases), and additional preference annotations between responses (Wang et al., 2024a).
-We adopt it as our retrieval corpus since (i) its richer annotations and multi-turn coverage yield more representative reasoning assets, (ii) preference signals provide finer-grained evidence for our Analyzer and Refiner agents, and (iii) it is the current benchmark standard for training and evaluating reward models, ensuring alignment with best practices in the field.
-Table 1. Overview of HelpSteer1/2 Annotation Dimensions. Each scored on a 0–4 scale.
-Attribute Description
-Helpfulness Overall helpfulness of the response to the prompt.
-Correctness Inclusion of all pertinent facts without errors.
-Coherence Consistency and clarity of expression.
-Complexity Intellectual depth required to write the response (e.g., basic competency vs. domain expertise).
-Verbosity Amount of detail included in the response relative to what is asked.
+[^1] The detailed prompts and specific role descriptions for agents are provided in Appendix H.
+4 Experimental design and setup
+4.1 Datasets
+We conduct experiments on the HelpSteer datasets (Wang et al., 2024c; b), which provide human-annotated prompt-response pairs scored on five quality dimensions using a 0–4 scale.
+[^2] Detailed descriptions of the five annotation dimensions are provided in Table 4 in Appendix A.
+Both datasets follow the same annotation schema but differ in size and construction, as summarized in Table 1.
+We build the retrieval corpus solely from the HelpSteer2 training set, which offers higher-quality annotations, broader prompt diversity, and preference labels between responses that provide richer supervision for our Analyzer and Refiner agents.
+HelpSteer2 also serves as a standard benchmark for reward modeling.
+Evaluation is conducted on the validation splits of HelpSteer1 and HelpSteer2.
+Dataset Train Val Multi-turn Preference Ann.
+HelpSteer1 35.3k 1.79k × ×
+HelpSteer2 20.3k 1.04k
+Table 1: Statistics of the HelpSteer1/2 datasets. Both share the same annotation schema of five quality dimensions scored on a 0–4 scale.
+4.2 Models used
 Reasoning–asset generation (corpus construction).
@@
 Generation models for evaluation.
-For downstream evaluation, we adopt a representative API model GPT-4o (Achiam et al., 2023), and an open-source model LLaMA-3-8B-Instruct (Dubey et al., 2024) as the backbone models.
+For downstream evaluation, we adopt a representative API model GPT-4o (Achiam et al., 2023), and an open-source model Llama3-8B-Instruct (Dubey et al., 2024) as the backbone models.
 These models are responsible for generating (i) optimized prompts $\hat{p}$ from the Refiner Agent and (ii) the corresponding optimized responses $\hat{r}$ conditioned on $\hat{p}$.
 Both models are configured with temperature =0=0 to ensure deterministic decoding, and all other hyperparameters remain at their defaults.
+Judge model.
+We employ the ArmoRM-Llama3-8B-v0.1 reward model (Wang et al., 2024a) as an interpretable multi-objective evaluator ($90\%$ benchmark accuracy), which automatically scores response quality.
+For any prompt-response pair, the judge model returns five HelpSteer-aligned scores: [^3] A detailed description of the ArmoRM model architecture, multi-objective training methodology, and alignment with HelpSteer dimensions is provided in Appendix C.
+$\mathcal{M}=\{\text{help},\text{corr},\text{coh},\text{comp},\text{verb}\},\quad s_{m}\in[0,4].$
 Retrieval.
 We adopt a sparse retriever based on BM25 (Robertson et al., 2009) over prompt text.
-For each test prompt $p_{\text{test}}$, we retrieve the top-$k$ training prompts with $k=3$, which we identify as the optimal setting for retrieval [^3] A detailed analysis of $k$ values and their effects on performance is in Section 5.1.1.
+For each test prompt $p_{\text{test}}$, we retrieve the top-$k$ training prompts with $k=3$, which we identify as the optimal setting for retrieval [^4] A detailed analysis of $k$ values and their effects on performance is provided in Section 5.2.1.
 and attach their paired responses and reasoning assets:
-$\{(p_{j},r_{j},\mathcal{R}_{j})\}_{j=1}^{k}.$
-4.2. Evaluation Method and Metrics
-Judge model.
-We employ the ArmoRM-Llama3-8B-v0.1 reward model (Wang et al., 2024b) as an interpretable multi-objective evaluator ($90\%$ benchmark accuracy), which automatically scores response quality.
-For any prompt-response pair, the judge model returns five HelpSteer-aligned scores (shown in Table 1):
-$\mathcal{M}=\{\text{help},\text{corr},\text{coh},\text{comp},\text{verb}\},\quad s_{m}\in[0,4].$
-Evaluation Metrics.
-Each candidate pair $(p,r)$ is evaluated across the five HelpSteer dimensions $\mathcal{M}=\{\text{help},\text{corr},\text{coh},\text{comp},\text{verb}\}$, where each raw score $s_{m}\in[0,4]$.
-We normalize scores by dividing by $4$, yielding $\tilde{s}_{m}=s_{m}/4\in[0,1]$.
-The overall quality score is then defined as the average over all metrics:
-$\mathrm{Score}(p,r)=\tfrac{1}{|\mathcal{M}|}\sum_{m\in\mathcal{M}}\tfrac{s_{m}}{4}.$
-This produces a single composite value in the range $[0,1]$, reflecting the overall response quality.
-4.3. Prompt Design
-To guide each agent in MA-SAPO, we carefully designed specialized prompts tailored to their roles in both the training and test phases.
-[^4] For brevity, we show shortened prompts; full templates are in our code repository.
-  1. (1)
-    The training prompts (Metric Explainer, Diagnostician, and Action Synthesizer) focus on generating structured reasoning assets by interpreting evaluation scores, diagnosing weaknesses, and synthesizing actionable improvements.
-  2. (2)
-    The test prompts (Analyzer and Refiner) leverage these assets to contrast retrieved exemplars with the original prompt and iteratively refine it while preserving the user’s intent.
-These prompts form the backbone of MA-SAPO’s reasoning-driven optimization pipeline, ensuring interpretability, controllability, and consistency across all stages.
-The prompt templates are provided below:
-Prompt Templates for Different Agents in Training and Testing Phases Metric Explainer Agent Prompt You are an evaluation explainer.
-Given a user prompt, a model response, and five 0–4 scores (helpfulness, correctness, coherence, complexity, verbosity), write ONE cohesive paragraph (6–9 sentences) that explains WHY each score was assigned and HOW to improve the response.
-Rules: - Output must be a single paragraph in plain text (no lists, no JSON, no headings) (…) Diagnostician Agent Prompt You are a precise evaluation doctor for LLM outputs.
-Task: Write ONE cohesive paragraph that (1) diagnoses the main Strength and Weaknesses, (2) explains why they matter in terms of the five metrics, and (3) prescribes concrete fixes.
-Rules: - Evidence-first: rely only on the prompt + response + reasoning.
-- Mention each metric once (helpfulness, correctness, coherence, complexity, verbosity) with a brief, specific rationale (…) Output: One paragraph.
-Action Synthesizer Agent Prompt Mission: Based on the structured diagnosis (Diagnostician Agent), provide a single actionable prompt suggestion that guides improvements while preserving the original intent.
-Focus on highlighting strengths, addressing weaknesses, and suggesting … Task: Using the diagnosis, propose concise guidance that: 1.
-Maintains the core topic and original goal of the prompt.
-2.
-Suggests ways to strengthen depth, clarity, and structure.
-3.
-Recommends addition or adjustment (…) Analyzer Agent Prompt You are the Analyzer Agent.
-Task: (1) Independently evaluate the original prompt (strengths and weaknesses).
-(2) Analyze retrieved prompt–response–feature triples from three agents.
-(3) Integrate insights into a unified analysis.
-Rules: - Begin with direct evaluation of the original prompt.
-- Evaluate each agent’s output concisely (strengths/weaknesses) (…) Output: A structured but cohesive paragraph-length analysis.
-Refiner Agent Prompt You are the Refiner Agent.
-Task: Using Analyzer feedback, optimize the original prompt to resolve all issues while preserving intent.
-Rules: - Extract the core improvement points from the Analyzer.
-- Prioritize Action Synthesizer suggestions over others (…) Output: Final optimized prompt.
-4.4. Baselines
-We implement six baselines under three different categories: single-pass without retrieval (1, 2, and 3), retrieval-augmented without reasoning assets (4), and multi-agent prompt optimization (5 and 6).
-All methods generate an optimized prompt from the initial prompt, and the final response is produced from that optimized prompt by the same backbone LLM used for the method.
-Specifically, the six baseline methods are detailed below:
-  1. (1)
-    Direct Generation: The LLM rewrites the input prompt into an optimized one, without intermediate reasoning.
-  2. (2)
-    Chain-of-Thought (CoT) [2022] (Wei et al., 2022): The LLM performs step-by-step reasoning about how to improve the prompt, then outputs an optimized version.
-  3. (3)
-    Role Assignment [2023] (Shanahan et al., 2023): A crafted system role (e.g., “You are a meticulous assistant optimizing prompts for helpfulness, correctness, coherence, complexity, and verbosity.”) guides the LLM to output an optimized prompt in one pass.
-  4. (4)
-    Retrieval-Augmented Generation (RAG) [2020] (Lewis et al., 2020): M25 retrieves the top-$k=10$ training prompts with their responses and corresponding scores.
-The LLM references these exemplars to produce an optimized prompt.
-  5. (5)
-    MAD [2023] (Du et al., 2023): MAD (Multi-Agent Debate) framework organizes several LLM “debaters” that each propose an answer with a rationale, then engage in multi-round critiques of one another’s reasoning.
-A separate judge (or a consensus rule) evaluates the arguments at each round and selects the final answer, which measurably improves factuality and step-by-step reasoning on strategic-reasoning tasks.
-  6. (6)
-    MARS [2025] (Zhang et al., 2025): MARS is a hierarchical system of seven agents.
-A Planner designs the optimization trajectory, while a Teacher, Critic, Student Socratic dialogue iteratively refines the prompt.
-At each iteration, a Target module evaluates the candidates, logs the history, and the system finally outputs the best prompt.
-Table 2.
-Main results on HelpSteer1 and HelpSteer2.
+$\{(p_{j},r_{j},\mathcal{R}_{j})\}_{j=1}^{k=3}.$ (6)
+4.3 Evaluation metrics
+We evaluate each candidate pair $(p,r)$ on the five HelpSteer dimensions $\mathcal{M}=\{\text{help},\text{corr},\text{coh},\text{comp},\text{verb}\}$, where each raw score $s_{m}\in[0,4]$.
+Scores are normalized as $\tilde{s}_{m}=s_{m}/4\in[0,1]$, and the overall quality score is computed as the average across all dimensions:
+$\mathrm{Score}(p,r)=\tfrac{1}{|\mathcal{M}|}\sum_{m\in\mathcal{M}}\tfrac{s_{m}}{4}.$ (7)
+This yields a single composite score in $[0,1]$ representing overall response quality.
+We also conduct a human evaluation with 17 expert annotators to assess (H1) the usefulness, accuracy, and consistency of the reasoning, and (H2) whether optimized prompts preserve the original intent.
+[^5] Human evaluation results are reported in Section 5.3, with detailed protocols and per-dimension scores in Appendix F.
+4.4 Baselines
+We compare our method with six baselines in three categories: single-pass prompt optimization without retrieval (Direct Generation, CoT (Wei et al., 2022), Role Assignment (Wu et al., 2023)), retrieval-augmented optimization (RAG (Lewis et al., 2020)), and multi-agent optimization (MAD (Du et al., 2024), MARS (Zhang et al., 2025)).
+For all methods, the model first generates an optimized prompt and then produces the final response using the same backbone LLM.
+We evaluate all baselines with GPT-4o and Llama3-8B-Instruct, using temperature 0 and default decoding settings.
+For RAG, we retrieve the top-$10$ training examples with BM25; for MAD, we use $2$ agents and $3$ debate rounds; and for MARS, we use 3344 planning steps and $5$ refinement iterations.
+[^6] Baseline details are provided in Appendix B.
+5 Experimental results and analysis
+5.1 Main results
+Methods GPT-4o Llama3-8B
+Help Corr Coh Comp Verb Avg Help Corr Coh Comp Verb Avg
+HelpSteer1
+Direct Generation 0.3216 0.3866 0.7583 0.2951 0.4215 0.4366 0.2549 0.3247 0.7054 0.2702 0.4062 0.3927
+RAG (sparse, k=10k{=}10) (2020) 0.3751 0.4402 0.7871 0.3116 0.4779 0.4784 0.2930 0.3594 0.7452 0.3007 0.4377 0.4272
+Chain-of-Thought (CoT) (2022) 0.3223 0.3876 0.7595 0.2935 0.4192 0.4364 0.2359 0.3077 0.6906 0.2627 0.3967 0.3787
+Role Assignment (2023) 0.3988 0.4679 0.8024 0.3427 0.5008 0.5025 0.2887 0.3516 0.7287 0.3164 0.4517 0.4274
+MAD (2024) 0.3774 0.4764 0.7954 0.3210 0.5248 0.4990 0.3774 0.5049 0.8067 0.4084 0.6708 0.5531
+MARS (2025) 0.4159 0.4876 0.7963 0.3293 0.5188 0.5096 0.3901 0.4591 0.7745 0.3512 0.5801 0.5110
+MA-SAPO (Ours) 0.5183 0.6260 0.8614 0.5013 0.7363 0.6486 0.4110 0.4868 0.8326 0.4720 0.8433 0.6091
+HelpSteer2
+Direct Generation 0.3616 0.4700 0.7723 0.3280 0.4913 0.4846 0.2616 0.3636 0.7078 0.2942 0.4421 0.4139
+RAG (sparse, k=10k{=}10) (2020) 0.4903 0.5745 0.8642 0.4161 0.6567 0.6003 0.3990 0.4711 0.7989 0.3722 0.5814 0.5245
+Chain-of-Thought (CoT) (2022) 0.2981 0.3958 0.7169 0.2888 0.4709 0.4341 0.1600 0.2545 0.6291 0.2421 0.3812 0.3334
+Role Assignment (2023) 0.4400 0.5175 0.8221 0.4025 0.5992 0.5563 0.2555 0.3303 0.7050 0.3267 0.4441 0.4123
+MAD (2024) 0.4167 0.5049 0.8067 0.4084 0.6708 0.5615 0.3971 0.4532 0.7898 0.3998 0.6439 0.5368
+MARS (2025) 0.4791 0.5569 0.8268 0.4095 0.6234 0.5791 0.4296 0.5019 0.7994 0.3957 0.6181 0.5482
+MA-SAPO (Ours) 0.5072 0.6038 0.8527 0.5244 0.7570 0.6490 0.4005 0.4754 0.8294 0.4833 0.8441 0.6065
+Table 2: Main results on HelpSteer1 and HelpSteer2.
 Columns report the five HelpSteer metrics (Help, Corr, Coh, Comp, Verb) normalized to $[0,1]$ and their mean (Avg).
 Methods are grouped into (i) single-pass prompting, (ii) retrieval-augmented generation (no reasoning assets), and (iii) multi-agent frameworks.
-Best within each backbone is bold; second best is underlined.
-Dataset Methods GPT-4o LLaMA-3-8B
+For each dataset and backbone block, the best value in each metric column is shown in bold, and the second-best is underlined.
+In our main experiments, we compare MA-SAPO with three groups of baselines: (1) single-pass prompting methods (Direct Generation, CoT, and Role Assignment), (2) retrieval-augmented generation (RAG) without reasoning assets, and (3) multi-agent frameworks (MAD and MARS).
+As shown in Table 2, MA-SAPO consistently achieves the best results across both HelpSteer1 and HelpSteer2 under both GPT-4o and Llama3-8B.
+Single-pass methods remain limited because they revise prompts without diagnostic grounding, often leading to brittle or inconsistent improvements.
+Role Assignment provides more explicit guidance, but still depends heavily on predefined roles.
+RAG benefits from retrieved exemplars, but without explicit reasoning about their usefulness, its refinements remain heuristic and less reliable.
+Figure 2: KDE score distributions of Direct Generation and MA-SAPO across five HelpSteer dimensions and their average, aggregated over GPT-4o and Llama3-8B.
+Solid lines denote HelpSteer1 and dashed lines denote HelpSteer2.
+Vertical lines indicate the mean of each distribution.
+MA-SAPO consistently shifts the score distribution toward higher values.
+Methods GPT-4o Llama3-8B
 Help Corr Coh Comp Verb Avg Help Corr Coh Comp Verb Avg
-HelpSteer1 Direct Generation 0.3216 0.3866 0.7583 0.2951 0.4215 0.4366 0.2549 0.3247 0.7054 0.2702 0.4062 0.3927
-Chain-of-Thought (CoT) 0.3223 0.3876 0.7595 0.2935 0.4192 0.4364 0.2359 0.3077 0.6906 0.2627 0.3967 0.3787
-Role Assignment 0.3988 0.4679 0.8024 0.3427 0.5008 0.5025 0.2887 0.3516 0.7287 0.3164 0.4517 0.4274
-RAG (sparse, k=10k{=}10) 0.3751 0.4402 0.7871 0.3116 0.4779 0.4784 0.2930 0.3594 0.7452 0.3007 0.4377 0.4272
-MAD 0.3774 0.4764 0.7954 0.3210 0.5248 0.4990 0.3774 0.5049 0.8067 0.4084 0.6708 0.5531
-MARS 0.4159 0.4876 0.7963 0.3293 0.5188 0.5096 0.3901 0.4591 0.7745 0.3512 0.5801 0.5110
-MA-SAPO (Ours) 0.5183 0.6260 0.8614 0.5013 0.7363 0.6486 0.4110 0.4868 0.8326 0.4720 0.8433 0.6091
-HelpSteer2 Direct Generation 0.3616 0.4700 0.7723 0.3280 0.4913 0.4846 0.2616 0.3636 0.7078 0.2942 0.4421 0.4139
-Chain-of-Thought (CoT) 0.2981 0.3958 0.7169 0.2888 0.4709 0.4341 0.1600 0.2545 0.6291 0.2421 0.3812 0.3334
-Role Assignment 0.4400 0.5175 0.8221 0.4025 0.5992 0.5563 0.2555 0.3303 0.7050 0.3267 0.4441 0.4123
-RAG (sparse, k=10k{=}10) 0.4903 0.5745 0.8642 0.4161 0.6567 0.6003 0.3990 0.4711 0.7989 0.3722 0.5814 0.5245
-MAD 0.4167 0.5049 0.8067 0.4084 0.6708 0.5615 0.3971 0.4532 0.7898 0.3998 0.6439 0.5368
-MARS 0.4791 0.5569 0.8268 0.4095 0.6234 0.5791 0.4296 0.5019 0.7994 0.3957 0.6181 0.5482
-MA-SAPO (Ours) 0.5072 0.6038 0.8527 0.5244 0.7570 0.6490 0.4005 0.4754 0.8294 0.4833 0.8441 0.6065
-For all baselines, we adopt GPT-4o and LLaMA-3-8B-Instruct as the backbone models, ensuring consistency and fairness in comparison by aligning with our main experimental setup.
-Baseline experiments were conducted with the following settings for fairness and fairness of our framework: all LLM temperatures were set to 0 and all other settings were set to LLM defaults.
-For MAD, it was fixed as $agent=2$ and $round=3$, and MARS designed the planner’s steps in steps 343-4 and set the entire iteration to $5$.
-Table 3.
-Ablation Studies on Retrieval Depth and Test Agent Combination.
-Varying the number of retrieved training examples ($k$) shows that $k=3$ (our default, denoted as MA-SAPO (k=3)) provides the best balance between contextual coverage and robustness.
-Collapsing the Analyzer and Refiner into a single agent yields competitive but less consistent results.
-Best is highlighted in bold; second best is underlined.
-Dataset Methods GPT-4o LLaMA-3-8B
-Help Corr Coh Comp Verb Avg Help Corr Coh Comp Verb Avg
-HelpSteer1 $k=1$ 0.5182 0.6247 0.8543 0.4943 0.7349 0.6453 0.4045 0.4769 0.8268 0.4729 0.8446 0.6051
+HelpSteer1
+$k=1$ 0.5182 0.6247 0.8543 0.4943 0.7349 0.6453 0.4045 0.4769 0.8268 0.4729 0.8446 0.6051
 $k=2$ 0.5211 0.6287 0.8591 0.4960 0.7318 0.6473 0.4002 0.4747 0.8236 0.4726 0.8384 0.6019
@@
 Test Agents Combination 0.4672 0.5415 0.8619 0.3670 0.6403 0.5756 0.4649 0.5203 0.8323 0.4363 0.7249 0.5957
-MA-SAPO (k=3) 0.5183 0.6260 0.8614 0.5013 0.7363 0.6486 0.4110 0.4868 0.8326 0.4720 0.8433 0.6091
-HelpSteer2 $k=1$ 0.5058 0.5982 0.8458 0.5113 0.7402 0.6403 0.3910 0.4616 0.8215 0.4784 0.8420 0.5989
+MA-SAPO (kk=3) 0.5183 0.6260 0.8614 0.5013 0.7363 0.6486 0.4110 0.4868 0.8326 0.4720 0.8433 0.6091
+HelpSteer2
+$k=1$ 0.5058 0.5982 0.8458 0.5113 0.7402 0.6403 0.3910 0.4616 0.8215 0.4784 0.8420 0.5989
 $k=2$ 0.5107 0.6044 0.8504 0.5158 0.7450 0.6452 0.3980 0.4713 0.8268 0.4845 0.8396 0.6041
@@
 Test Agents Combination 0.5404 0.6012 0.8891 0.4455 0.7398 0.6432 0.4379 0.4720 0.7968 0.4918 0.7321 0.5861
-MA-SAPO (k=3) 0.5072 0.6038 0.8527 0.5244 0.7570 0.6490 0.4005 0.4754 0.8294 0.4833 0.8441 0.6065
-5. Experimental Results and Analysis
-5.1. Main Results
-In our main experiments, we compare MA-SAPO with three families of baselines: (1) single-pass prompting methods including Direct Generation, Chain-of-Thought (CoT), and Role Assignment, (2) retrieval-augmented generation (RAG) without reasoning assets, and (3) multi-agent frameworks such as MAD and MARS.
-As outlined in Section 4.2, we evaluate on both HelpSteer1 and HelpSteer2 using the five annotated quality dimensions and their normalized average score as metrics.
-The results are shown in Table 2.
-In detail, Single-pass methods provide the simplest form of optimization, but their weakness lies in performing edits in isolation.
-Direct Generation and CoT are essentially one-shot rewrites that lack diagnostic grounding, which makes their improvements brittle and often inconsistent across metrics.
-Role Assignment can guide the model to consider multiple aspects explicitly, yet it still suffers from over-reliance on handcrafted role descriptions and cannot adapt when those roles fail to cover task-specific subtleties.
-RAG improves upon these by injecting exemplars, but the optimization process remains unguided: retrieved examples are consumed without explanation of why they are useful, leaving the refinement process heuristic and prone to inconsistency.
-As a result, these methods either underperform on complex prompts or achieve higher scores only at the cost of verbosity and drift in task semantics.
-Recent multi-agent frameworks such as MAD and MARS perform better by coordinating multiple roles in a structured workflow, yet they remain outcome-driven and heavy in cost [^5] A detailed cost and latency comparisons are provided in Section 5.4..
-Their reliance on repeated debate or critique cycles incurs significant computational overhead, while still treating evaluation as a scalar target rather than a source of actionable reasoning.
-In contrast, our MA-SAPO consistently surpasses all baselines across datasets and backbones by explicitly converting evaluation signals into reusable reasoning assets.
-The Analyzer to Refiner pipeline applies only evidence-backed edits drawn from these assets, ensuring that optimizations are not only effective but also interpretable and controllable [^6] A detailed case study is provided in Section 5.5..
-For instance, on HelpSteer1 with GPT-4o, MA-SAPO achieves 0.6486 average score compared to 0.5096 for MARS and 0.4784 for RAG.
-Similar margins hold across HelpSteer2 and with LLaMA-3-8B, confirming that structured reasoning, rather than trial-and-error search, is the main driver of performance.
-Overall, MA-SAPO combines accuracy and efficiency, making it a more practical and principled framework for prompt optimization than competitive baselines.
-5.2. Ablation Study
-5.2.1. Top-k Retrieval Variations
-To examine how the number of retrieved training examples influences the performance of MA-SAPO, we vary the retrieval depth $k\in\{1,2,3,4\}$.
-Table 3 reports results for both GPT-4o and LLaMA-3-8B models across HelpSteer1 and HelpSteer2.
-For both backbones, increasing $k$ from 1 to 3 generally improves overall performance.
-With $k=1$, the Analyzer is under-informed, as it can only compare the test prompt against a single exemplar, limiting the diversity of reasoning assets available.
-Moving to $k=2$ provides broader evidence and modest gains, while $k=3$ yields the best balance across all five metrics and achieves the highest average performance.
-At $k=4$, results remain competitive but show slight declines in some dimensions, suggesting that introducing too many exemplars adds noise and creates conflicting diagnostic signals.
-These findings highlight a clear trade-off: smaller $k$ restricts contextual diversity, while larger $k$ risks diluting the relevance of retrieved assets.
-Overall, $k=3$ achieves the most consistent and robust performance across datasets and backbones, and we adopt it as the default retrieval setting in all subsequent experiments.
-5.2.2. Test Agents Combination
-To further assess the robustness of MA-SAPO, we conduct an ablation study where the two test-phase agents are merged into a single combined agent, referred to as Test Agents Combination.
-Table 3 summarizes the results for both GPT-4o and LLaMA-3-8B on HelpSteer1 and HelpSteer2.
-We observe that the combined variant achieves competitive performance, occasionally surpassing the full MA-SAPO in a single metric.
-However, the average performance consistently falls behind the full framework.
-This suggests that while a unified agent can capture certain improvements, it lacks the structured diagnostic pipeline of MA-SAPO, where the Analyzer first produces evidence-grounded reports and the Refiner then executes targeted edits.
-These findings demonstrate the necessity of separating the two roles in the test phase.
-The Analyzer to Refiner pipeline enforces a more interpretable and controllable workflow, ensuring that refinements are justified by retrieved reasoning assets rather than introduced implicitly by a single agent.
-As a result, the modular design of MA-SAPO is not only more reliable but also essential for achieving consistent gains across datasets and backbones.
-5.3. Qualitative Analysis
-To complement the automatic benchmark results, we conducted a human evaluation designed to assess both the interpretability and controllability of MA-SAPO’s optimization process.
-This evaluation focused on two complementary hypotheses: (H1) that multi-agent reasoning produces higher-quality reasoning than a single-agent baseline, and (H2) that MA-SAPO preserves the semantic direction and intent of the original prompts during optimization.
-All evaluations were conducted by 14 independent human annotators (mean age = 29.1, SD = 3.4), who were either data scientists, AI engineers, or graduate students majoring in data science or related fields.
+MA-SAPO (kk=3) 0.5072 0.6038 0.8527 0.5244 0.7570 0.6490 0.4005 0.4754 0.8294 0.4833 0.8441 0.6065
+Table 3: Ablation results over agent number $k$ and agent-combination testing on HelpSteer1 and HelpSteer2. Best scores are shown in bold and second-best scores are underlined.
+Multi-agent baselines such as MAD and MARS achieve stronger performance by coordinating multiple roles, but they still rely on repeated interaction cycles and treat evaluation mainly as a scalar target.
+In contrast, MA-SAPO converts evaluation outcomes into reusable reasoning assets and uses them to support evidence-backed refinement through the Analyzer-to-Refiner pipeline, yielding stronger and more interpretable improvements.
+For example, on HelpSteer1 with GPT-4o, MA-SAPO achieves an average score of 0.6486, outperforming MARS (0.5096) and RAG (0.4784).
+Similar trends hold across HelpSteer2 and Llama3-8B, and Figure 2 further shows that these gains reflect a consistent shift toward higher scores across dimensions rather than a few outlier cases.
+[^7] Per-backbone KDE distributions are provided in Appendix E.
+This suggests that MA-SAPO improves multiple quality aspects simultaneously, rather than over-optimizing a single metric.
+5.2 Ablation study
+5.2.1 Top-k retrieval variations
+To further analyze the robustness of MA-SAPO, we examine two key design choices: retrieval depth and test-phase agent composition.
+First, to study how the number of retrieved training examples affects performance, we vary the retrieval depth $k\in\{1,2,3,4\}$.
+Table 3 reports results for both GPT-4o and Llama3-8B across HelpSteer1 and HelpSteer2.
+Across both backbones, increasing $k$ from 1 to 3 generally improves performance, with $k=3$ achieving the best average balance across the five metrics.
+When $k=1$, the Analyzer is limited to a single exemplar, restricting the diversity of available reasoning assets.
+Increasing $k$ broadens the evidence base, but $k=4$ shows slight declines in some dimensions, suggesting that retrieving too many exemplars can introduce noise and conflicting signals.
+Overall, these results indicate a clear trade-off between contextual diversity and diagnostic noise, with $k=3$ providing the most reliable setting in our experiments.
+5.2.2 Test agents combination
+We subsequently assess the effect of test-phase agent modularity by merging the two agents into a single combined agent, denoted as Test Agents Combination.
+As shown in Table 3, this variant remains competitive for both GPT-4o and Llama3-8B on HelpSteer1 and HelpSteer2, and occasionally surpasses the full framework on individual metrics.
+However, its average performance consistently falls below that of the full MA-SAPO, indicating that it cannot reliably reproduce the benefits of the structured Analyzer-to-Refiner workflow.
+By separating evidence-grounded diagnosis from targeted refinement, the modular design supports a more interpretable and controllable process and leads to more consistent gains across datasets and model backbones.
+5.3 Qualitative analysis
+We additionally conducted a human evaluation to examine two complementary aspects of MA-SAPO.
+[^8] The full evaluation protocol, per-dimension scores, and the rating distribution are provided in Appendix F.
+A total of 17 annotators, comprising data scientists, AI engineers, and graduate students in data science or related fields, participated in the study.
+For reasoning quality (H1), MA-SAPO achieved a mean score of 3.93 on a 5-point scale, outperforming the single-agent baseline by 0.24 points, with notable gains in usefulness and accuracy.
+For directional consistency (H2), annotators assessed whether optimized prompts retained the original intent, yielding a mean rating of 3.36 out of 4.0, indicating that prompt optimization largely preserves semantic alignment.
+Taken together, the human evaluation supports the quantitative findings and demonstrates that MA-SAPO improves output quality without distorting the original task intent.
+5.4 Cost and efficiency analysis
+Figure 3: Efficiency–performance trade-off of MAD, MARS, and MA-SAPO.
+Panel (a) plots average input tokens per instance against average score, and panel (b) plots execution time per instance against average score, under GPT-4o and Llama3-8B.
+Averaged over HelpSteer1 and HelpSteer2, MA-SAPO delivers the highest scores with the lowest token cost and substantially lower latency than MARS.
+To complement the effectiveness results, we also compare the computational cost of different multi-agent frameworks.
+Figure 3(a) reports the average input tokens per instance against average score.
+MAD incurs moderate overhead but yields relatively weak overall performance, while MARS improves performance at a substantially higher token cost, consuming over 17K input tokens per instance.
+In contrast, MA-SAPO reduces token consumption to approximately 5K input tokens per instance by shifting reasoning asset construction offline and using a single Analyzer-to-Refiner loop at test time, while achieving the highest average scores under both GPT-4o and Llama3-8B.
+Figure 3(b) further illustrates the trade-off between execution time per instance and average score.
+MAD achieves the shortest execution time at approximately 5 seconds per instance, yet delivers the lowest average performance among the three methods.
+MARS incurs the highest latency, requiring roughly 85 to 90 seconds per instance.
+By contrast, MA-SAPO maintains an execution time of approximately 40 seconds per instance, which is less than half that of MARS, while delivering substantially higher scores.
+Taken together, these results show that MA-SAPO achieves the most favorable efficiency-performance balance across both token cost and latency.
+[^9] Detailed model inference hardware specifications are provided in Appendix D.
+6 Conclusion
+We propose MA-SAPO, a novel multi-agent framework for interpretable and controllable prompt optimization.
+Compared to prior methods that either perform one-shot rewrites, rely on unguided retrieval, or employ heavy multi-agent debates, MA-SAPO explicitly transforms evaluation outcomes into reusable reasoning assets and grounds refinements in evidence-backed Analyzer-to-Refiner interactions.
+[^10] Practical implications are discussed in Appendix G.
+Our contributions extend beyond framework design to include comprehensive experiments on two benchmarks and multiple backbones, ablation studies on retrieval depth and agent configuration, and human evaluations of reasoning quality and semantic consistency.
+Extensive results demonstrate that MA-SAPO consistently outperforms existing approaches, achieving higher optimization quality while remaining efficient in cost and latency.
+Limitations
+Although MA-SAPO achieves superior performance compared to a range of single-agent reasoning baselines and interaction-based agent frameworks, these results do not imply that MA-SAPO uniformly dominates all baselines across tasks and conditions.
+MA-SAPO leverages retrieved exemplars as evidence for diagnosis and prompt rewriting.
+However, retrieval may suffer from limited recall under paraphrase or lexical mismatch, causing the system to rely on less relevant assets and creating a bottleneck at the retrieval stage.
+Moreover, the single-path, sequential multi-agent pipeline is prone to cascading errors and provides limited exploration, since it does not explicitly branch over, compare, and select among alternative diagnostic/edit trajectories.
+Future work could improve robustness by adding verifier-based artifact validation, uncertainty-aware gated updates, and constrained rewriting, as well as by introducing multi-path candidate generation with iterative verification.
+Appendix
+Appendix A Datasets detail
+We conduct our experiments on the HelpSteer family of datasets (Wang et al., 2024c; b), which provide human-annotated prompt-response pairs for evaluating prompt optimization in large language models.
+Both datasets share the same annotation schema across five quality dimensions, but differ in size, design, and annotation depth, as summarized in Table 4.
+We build our retrieval corpus from the HelpSteer2 training set and conduct evaluation on the validation splits of both datasets.
+Attribute Description
+Helpfulness Overall helpfulness of the response to the prompt.
+Correctness Inclusion of all pertinent facts without errors.
+Coherence Consistency and clarity of expression.
+Complexity Intellectual depth required to write the response.
+Verbosity Amount of detail included in the response relative to what is asked.
+Table 4: Overview of HelpSteer1/2 annotation dimensions. Each scored on a 0–4 scale.
+Figure 4: Score distributions across five annotation dimensions for HelpSteer1 (left) and HelpSteer2 (right).
+Each bar shows the proportion of samples assigned each score from 0 to 4.
+Coherence is heavily concentrated at higher scores in both datasets, while complexity and verbosity exhibit broader distributions across the full range.
+A.1 HelpSteer1
+HelpSteer1 contains 35.3k training and 1.79k validation samples, each comprising a prompt, a response, and five human-annotated attributes scored on a 0–4 scale.
+Prompts cover a broad range of task types including rewriting, summarization, classification, extraction, and question answering, with up to four responses generated per prompt.
+Each response was independently reviewed by multiple annotators, providing reliable quality estimates across diverse prompt-response styles.
+A.2 HelpSteer2
+HelpSteer2 is a newer dataset containing 20.3k training and 1.04k validation samples, with higher-quality annotations, multi-turn prompts (about 29% of cases), and additional preference annotations between responses (Wang et al., 2024b).
+Compared to HelpSteer1, it employs a substantially larger annotator pool and a more diverse set of response sources, with each sample receiving a minimum of three independent annotations.
+We adopt it as our retrieval corpus since (i) its richer annotations and multi-turn coverage yield more representative reasoning assets, (ii) preference signals provide finer-grained evidence for our Analyzer and Refiner agents, and (iii) it is the current benchmark standard for training and evaluating reward models, ensuring alignment with best practices in the field.
+As shown in Figure 4, HelpSteer2 exhibits more concentrated distributions toward higher scores, particularly in coherence, reflecting the higher quality of responses.
+Appendix B Baseline details
+We compare our method against six baselines from three categories: single-pass prompting, retrieval-augmented generation, and multi-agent frameworks.
+For all methods, the model first produces an optimized prompt from the initial prompt, and the final response is then generated from the optimized prompt using the same backbone LLM.
+B.1 Single-pass prompting
+This category includes Direct Generation, Chain-of-Thought (CoT) (Wei et al., 2022), and Role Assignment (Shanahan et al., 2023).
+These methods optimize the prompt without retrieving external examples.
+Direct Generation rewrites the input prompt into an optimized prompt in a single pass.
+Chain-of-Thought (CoT) generates intermediate reasoning steps before producing the optimized prompt.
+Role Assignment prepends a predefined role or persona to the input prompt and generates the optimized prompt in a single pass.
+B.2 Retrieval-augmented generation
+This category includes RAG (Lewis et al., 2020).
+The model is provided with the top-$k=10$ retrieved training examples as references when generating the optimized prompt.
+No additional reasoning assets are used.
+B.3 Multi-agent framework
+This category includes MAD (Du et al., 2024) and MARS (Zhang et al., 2025).
+MAD uses multiple agents to propose candidate prompt revisions, exchange feedback over several rounds, and select the final optimized prompt.
+MARS uses a hierarchical multi-agent framework that first plans the optimization procedure and then iteratively refines the prompt.
+For all baselines, we use GPT-4o and Llama3-8B-Instruct as the backbone models.
+We set the decoding temperature to 0 and keep all other generation settings at their default values.
+For MAD, we use $2$ agents and $3$ debate rounds.
+For MARS, the planner produces $3$ to $4$ planning steps, followed by $5$ refinement iterations.
+Appendix C LLM-as-a-judge
+To automatically evaluate the quality of responses generated from our HelpSteer-based prompt generation framework, we use ArmoRM-Llama3-8B-v0.1 as the LLM-as-a-judge.
+ArmoRM-Llama3-8B-v0.1 (Wang et al., 2024a) is a Llama-3 8B-based reward model that departs from conventional scalar preference modeling by decomposing response quality into multiple interpretable reward objectives.
+Specifically, the model first learns attribute-level rewards from multi-dimensional absolute-rating data through an Absolute-Rating Multi-Objective Reward Model (ArmoRM), and then aggregates them into a final preference score using a Mixture-of-Experts (MoE) gating network conditioned on the prompt context.
+This model is particularly suitable for our evaluation setting because it directly incorporates the five HelpSteer annotation dimensions (helpfulness, correctness, coherence, complexity, and verbosity) as internal reward objectives.
+Since the model is trained to predict these HelpSteer-related attributes as part of its multi-objective reward formulation, it is well aligned with our task of evaluating prompt-response pairs derived from a HelpSteer-based framework.
+In addition, its two-stage design separates attribute-level reward prediction from final scalar aggregation, which provides a more interpretable evaluation process than conventional single-score preference models.
+Empirically, ArmoRM-Llama3-8B-v0.1 achieves an overall score of 89.0 on RewardBench, outperforming the corresponding Llama-3 8B Bradley–Terry reward model (83.6).
+These reported results support our use of ArmoRM-Llama3-8B-v0.1 not only as an open and efficient judge model, but also as a reward model whose internal objectives are structurally aligned with the HelpSteer evaluation dimensions.
+Appendix D Hardware Specifications
+Model Hardware Power (TDP) Access
+Meta-Llama-3-8B-Instruct NVIDIA H100-80GB 700W Local GPU
+GPT-4o OpenAI API
+Table 5: Hardware specifications used for model inference.
+Table 5 summarizes the hardware configurations used for model inference in our experiments.
+For the open-source model, inference with Meta-Llama-3-8B-Instruct was conducted on NVIDIA H100-80GB GPU.
+For the proprietary model, GPT-4o was accessed via the OpenAI API, and no local GPU resources were required.
+Appendix E Per-backbone score distributions
+Figures 5 and 6 decompose the aggregated KDE distributions from Figure 2 by backbone model, reporting score distributions for GPT-4o and Llama3-8B separately.
+In each figure, the five HelpSteer dimensions and their average are shown as individual subplots, comparing Direct Generation against MA-SAPO across both HelpSteer1 and HelpSteer2.
+E.1 GPT-4o
+Figure 5: KDE score distributions for GPT-4o. Solid lines denote HelpSteer1 and dashed lines denote HelpSteer2. Vertical lines indicate the mean of each distribution.
+Under GPT-4o, MA-SAPO produces pronounced rightward shifts across all five dimensions.
+Helpfulness and complexity show the largest gains, with Direct Generation distributions concentrated around scores of 1.2–1.3 and MA-SAPO shifting the mass to 2.1–2.2.
+Verbosity exhibits the most dramatic absolute shift, with Direct Generation peaking near 1.9 and MA-SAPO reaching approximately 3.0.
+Correctness also improves substantially, with the distribution peak moving from roughly 1.8 to 2.5 while becoming more tightly concentrated.
+Coherence is the exception, as Direct Generation already achieves high scores near 3.0, leaving limited headroom for further improvement.
+Nevertheless, MA-SAPO still achieves a visible upward shift in the distribution mean.
+The Average subplot confirms that these per-metric gains compound into a consistent overall improvement across both HelpSteer1 and HelpSteer2.
+E.2 Llama3-8B
+Figure 6: KDE score distributions for Llama3-8B. Solid lines denote HelpSteer1 and dashed lines denote HelpSteer2. Vertical lines indicate the mean of each distribution.
+Under Llama3-8B, the directional pattern is consistent with GPT-4o, though with notable differences in distribution shape.
+Correctness under Direct Generation exhibits a broad, flat distribution across low scores, whereas MA-SAPO produces a sharp, concentrated peak near 2.0, indicating not only higher scores but also more consistent output quality.
+Verbosity shows the most dramatic shift among all metrics, with the MA-SAPO distribution peak reaching approximately 3.4 compared to roughly 1.8 for Direct Generation.
+Helpfulness and complexity also shift rightward, though with somewhat smaller absolute gains than under GPT-4o.
+As with GPT-4o, coherence shows the most modest improvement, consistent with its already high baseline scores.
+E.3 Cross-backbone comparison
+Across both backbones, MA-SAPO consistently shifts score distributions toward higher values in all five dimensions.
+Verbosity, complexity, helpfulness, and correctness show the largest shifts, while coherence improves more modestly due to the high baseline scores already achieved by Direct Generation.
+One notable difference between backbones is that Llama3-8B shows a sharper distributional improvement in correctness, with Direct Generation producing a much flatter spread compared to GPT-4o.
+Overall, the per-backbone breakdown confirms that the improvements observed in the aggregated figure are consistent across model families.
+Appendix F Qualitative analysis details
+This appendix provides the full details of the human evaluation summarized in Section 5.3.
+The evaluation focused on two complementary hypotheses: (H1) that multi-agent reasoning produces higher-quality reasoning than a single-agent baseline, and (H2) that MA-SAPO preserves the semantic direction and intent of the original prompts during optimization.
+All evaluations were conducted by 17 independent human annotators (mean age = 29.1, SD = 3.4), who were either data scientists, AI engineers, or graduate students majoring in data science or related fields.
 Each annotator was familiar with language model evaluation and prompt optimization procedures.
 The evaluation employed a 5-point Likert scale for reasoning quality (H1) and a 4-point scale for directional consistency (H2).
-The following subsections describe the evaluation setup and results.
-5.3.1. Evaluating Multi-Agent Reasoning Quality (H1)
+F.1 Evaluating multi-agent reasoning quality (H1)
 To assess whether MA-SAPO enhances reasoning quality, we randomly sampled 30 reasoning cases from the HelpSteer1 training set.
 Each case included two reasoning outputs: one generated by a single-agent baseline and one by MA-SAPO’s multi-agent reasoning pipeline, resulting in 60 reasoning outputs in total.
-These outputs were independently evaluated by 14 human annotators, yielding 840 individual evaluations (30 cases $\times$ 2 outputs $\times$ 14 annotators).
+These outputs were independently evaluated by 17 human annotators, yielding 1020 individual evaluations (30 cases $\times$ 2 outputs $\times$ 17 annotators).
 Each output was rated along three qualitative dimensions:
@@
 As summarized in Table 6, MA-SAPO outperformed the single-agent baseline across all three dimensions.
-The largest gains were observed in perceived usefulness and factual accuracy, both showing statistically significant improvements (*p* ¡ 0.05, paired $t$-test, $n=30$).
+The largest gains were observed in perceived usefulness and factual accuracy, both showing statistically significant improvements ($p<0.05$, paired $t$-test, $n=30$).
 These findings indicate that multi-agent reasoning yields more helpful, accurate, and consistent responses.
-Table 4.
-Human evaluation of reasoning quality.
-MA-SAPO outperforms the single-agent baseline in usefulness, factual accuracy, and consistency.
-Scores are averaged on a 1–5 scale; asterisks denote significance levels from paired $t$-tests ($^{*}p<0.05$, $n=30$).
 Method Usefulness Accuracy Consistency Mean
@@
 MA-SAPO (Ours) 3.89 3.87 4.02 3.93
-$\Delta$ (Improvement) +0.25 +0.24 +0.21 +0.23
-5.3.2. Evaluating Directional Consistency of Optimized Prompts (H2)
+$\Delta$ (Improvement) +0.25 +0.24 +0.21 +0.24
+Table 6: Human evaluation of reasoning quality.
+MA-SAPO outperforms the single-agent baseline in usefulness, factual accuracy, and consistency.
+Scores are averaged on a 1–5 scale; asterisks denote significance levels from paired $t$-tests ($^{*}p<0.05$, $n=30$).
+F.2 Evaluating directional consistency of optimized prompts (H2)
 To evaluate whether MA-SAPO preserves the semantic intent of prompts during optimization, we randomly sampled 40 prompt pairs, each consisting of an original prompt ($A$) and its optimized version ($B$) generated by MA-SAPO.
-Each pair was independently rated by 14 human annotators, resulting in a total of 560 evaluations (40 pairs $\times$ 14 annotators).
+Each pair was independently rated by 17 human annotators, resulting in a total of 680 evaluations (40 pairs $\times$ 17 annotators).
 Annotators rated the directional consistency of each pair on a 4-point Likert scale (1 = completely changed, 2 = partially changed, 3 = mostly preserved, 4 = fully preserved), indicating how well the optimized prompt maintained the intent, goal, and domain of the original.
@@
 Most scores are concentrated between 3 and 4, indicating that the optimized prompts largely preserved the original semantic intent.
-The mean rating was 3.36 with a standard deviation of 0.74 ($n=560$), demonstrating that MA-SAPO effectively enhances prompt clarity and structure while maintaining semantic stability.
+The mean rating was 3.36 with a standard deviation of 0.74 ($n=680$), demonstrating that MA-SAPO effectively enhances prompt clarity and structure while maintaining semantic stability.
 The relatively moderate standard deviation further suggests that annotators consistently judged the optimized prompts as semantically aligned with their originals.
-Figure 3.
-Distribution of directional consistency ratings (n=560).
+Figure 7: Distribution of directional consistency ratings (n=680).
 Ratings on a 4-point scale (1 = changed, 4 = preserved).
 The dashed line marks the mean (3.36) and the shaded band shows one standard deviation (SD = 0.74).
-5.3.3. Summary of Human Evaluation
+F.3 Summary of human evaluation
 The human evaluation results highlight two complementary aspects of MA-SAPO.
@@
 Second, MA-SAPO maintains strong semantic stability during prompt optimization, ensuring that improved clarity and structure do not distort the original task intent.
-Together, these findings demonstrate that MA-SAPO achieves both reasoning quality and semantic controllability, reinforcing the quantitative evidence presented in Section 5.3.
-5.4. Cost and Latency
-Table 5.
-Cost and Latency Comparison.
-MA-SAPO requires far fewer tokens and calls than other competitive multi-agent baselines, while maintaining superior optimization performance.
-Lower is better ($\downarrow$).
-Method # Input Tokens $\downarrow$ # Output Tokens $\downarrow$ # Calls $\downarrow$
-MAD 5,217 2,170 8
-MARS 17,149 6,789 60
-MA-SAPO (Ours) 4,968 526 2
-To complement the effectiveness results, we also compare the computational cost of different multi-agent frameworks.
-Table 5 reports the average number of input tokens, output tokens, and calls per instance.
-We observe that existing multi-agent systems such as MAD and MARS incur substantial overhead.
-MAD requires around 5K input tokens, 2K output tokens, and 8 calls on average due to repeated multi-round debates.
-MARS is even heavier, consuming more than 17K input tokens, nearly 7K output tokens, and about 60 calls per instance, reflecting its iterative planner–critic–student workflow.
-In contrast, MA-SAPO is significantly more efficient: by constructing reasoning assets offline and adopting a single Analyzer to Refiner loop at test time, it reduces the runtime cost to roughly 5K input tokens, 0.5K output tokens, and only 2 calls per instance.
-These results highlight that MA-SAPO not only improves optimization quality but also achieves a more favorable cost–latency trade-off.
-While MAD and MARS achieve moderate gains at the expense of heavy resource usage, MA-SAPO provides consistent improvements while keeping computational demands.
-Figure 4.
-Case study illustrating MA-SAPO’s multi-agent prompt optimization process.
-Starting from an initial user prompt and evaluation scores, the framework retrieves reasoning assets, analyzes weaknesses, and iteratively refines the prompt.
-The regenerated response shows improved clarity, structure, and contextual depth compared to the original.
-5.5. Case Study
-To further illustrate the effectiveness and practical utility of MA-SAPO, we present a case study drawn from the HelpSteer2 validation dataset.
+Together, these findings demonstrate that MA-SAPO achieves both reasoning quality and semantic controllability, reinforcing the quantitative evidence presented in the main text.
+Appendix G Implications
+This work shows that prompt optimization can move beyond black-box score chasing by explicitly converting evaluation outcomes into reusable reasoning assets (explanations, diagnoses, and edit directives) that support targeted, evidence-backed edits.
+In practice, MA-SAPO offers a more interpretable and controllable workflow: rather than relying on trial-and-error rewriting, practitioners can trace changes to retrieved exemplars and their metric-linked rationales, which is valuable when prompt updates must be reviewed, audited, or kept consistent across versions.
+Moreover, MA-SAPO enables cost-aware scaling by shifting most computation offline: once the reasoning-asset corpus is built, new prompts can be optimized with lightweight retrieval plus an Analyzer-to-Refiner step, making it suitable for maintaining prompt libraries in real systems (e.g., assistants, support bots, enterprise applications) while reducing expensive multi-round agent interactions.
+For instance, the key contribution of MA-SAPO is a contrastive mechanism that compares, verifies, and selects among multiple LLM-generated candidates, rather than relying on advanced retrieval.
+As a result, even with BM25 alone, our framework remains competitive while being cheaper and faster to run.
+Consequently, MA-SAPO turns evaluation traces into retrievable, auditable reasoning assets, enabling repeatable prompt updates with predictable cost.
+Appendix H Prompt template & design
+In this section, we provide the complete prompt templates used to orchestrate the collaborative reasoning of each agent within the MA-SAPO framework.
+As briefly discussed in the main text, these specialized instructions are categorized into two distinct phases.
+The training phase prompts guide the Metric Explainer, Diagnostician, and Action Synthesizer agents in generating structured reasoning assets from evaluation scores.
+Conversely, the test phase prompts direct the Analyzer and Refiner agents to sequentially optimize the user’s initial prompt based on the retrieved assets.
+The detailed role descriptions, core tasks, and strict output constraints for each agent are presented below.
+H.1 Training phase prompts
+Prompt template: Metric explainer agent
+Role: You are an evaluation explainer.
+Task: Explain WHY each of the five scores (0–4) was assigned to the model response.
+Provide a cohesive paragraph that bridges the gap between numerical scores and actionable improvements.
+Inputs: [USER_PROMPT]: {prompt_text} [MODEL_RESPONSE]: {model_response} [SCORES]: {5 metrics (helpfulness, correctness, coherence, complexity, verbosity)} Strict Rules: Output must be a single paragraph in plain text.
+Cover metrics in order: helpfulness $\rightarrow$ correctness $\rightarrow$ coherence $\rightarrow$ complexity $\rightarrow$ verbosity.
+Ground every rationale specifically in the provided response evidence.
+Maintain a neutral and specific tone.
+Table 7: Prompt template for the metric explainer agent.
+Prompt template: Diagnostician agent
+Role: You are a precise evaluation doctor for LLM outputs.
+Task: Diagnose the primary strengths and weaknesses of the response based on metric explanations.
+Prescribe concrete fixes to resolve identified reasoning or alignment gaps.
+Inputs: [USER_PROMPT]: {prompt_text} [MODEL_RESPONSE]: {model_response} [SCORES]: {5 metrics (helpfulness, correctness, coherence, complexity, verbosity)} [REASONING_PARAGRAPH]: {reasoning_paragraph} Strict Rules: Rely strictly on the prompt, response, and provided reasoning.
+Address each metric once with a brief, specific rationale.
+Flag any mismatches between the numerical scores and the qualitative reasoning.
+Conclude with a crisp prescription for the required response elements.
+Table 8: Prompt template for the diagnostician agent.
+Prompt template: Action synthesizer agent
+Role: You are a suggestion provider.
+Task: Synthesize the diagnosis into a single actionable prompt suggestion.
+Focus on enhancing the prompt’s structural and contextual depth without altering the original intent.
+Inputs: [USER_PROMPT]: {prompt_text} [MODEL_RESPONSE]: {model_response} [SCORES]: {5 metrics (helpfulness, correctness, coherence, complexity, verbosity)} [DIAGNOSIS]: {diagnosis_text} Strict Rules: Output ONLY the improvement suggestion (no conversational filler).
+Maintain the core topic and original objective of the user prompt.
+Focus on actionable directions (e.g., adding constraints, requesting specific formats).
+Balance the five metrics to guide the model toward a higher-quality response.
+Table 9: Prompt template for the action synthesizer agent.
+H.2 Test phase prompts
+Prompt template: Analyzer agent
+Role: You are the analyzer agent.
+Task: Conduct an independent evaluation of the prompt’s inherent qualities.
+Integrate insights from the three previous agents (Metric Explainer, Diagnostician, Action Synthesizer) into a unified analysis.
+Inputs: [USER_PROMPT]: {prompt} [RETRIEVED_CONTENT]: {retrieved_content} (Aggregated agent outputs) Strict Rules: Prioritize the Action Synthesizer’s output when synthesizing final improvement points.
+Eliminate redundant observations across agent outputs.
+Ensure all proposed improvements are concrete and actionable.
+Output the final analysis in a structured, cohesive paragraph.
+Table 10: Prompt template for the analyzer agent.
+Prompt template: Refiner agent
+Role: You are the refiner agent.
+Task: Optimize the original prompt based on the Analyzer Agent’s feedback.
+Produce a final, high-performance prompt that maximizes clarity and task effectiveness.
+Inputs: [USER_PROMPT]: {prompt} [ANALYZER_FEEDBACK]: {analyzed_result} Strict Rules: Directly incorporate the prioritized improvements from the feedback.
+Preserve the original intent and task objective perfectly.
+Remove any unnecessary elaboration to ensure the prompt remains concise.
+Output ONLY the final optimized prompt.
+Table 11: Prompt template for the refiner agent.
+Appendix I Case study
+To further illustrate the effectiveness and practical utility of MA-SAPO, we present three case studies drawn from the HelpSteer2 validation dataset.
 [^11] Additional case studies demonstrating diverse application scenarios are available in our code repository.
-This example highlights MA-SAPO’s strengths in collaborative reasoning, context-aware prompt refinement, and robust adaptation across diverse input types, thereby providing concrete evidence of the framework’s ability to enhance prompt optimization in complex reasoning scenarios.
-The case shown in Figure 4 demonstrates the end-to-end prompt optimization process.
-It begins with an initial input prompt, “What are 10 primary ways ChatGPT will change 4th–12th grade education in the United States?”, alongside the corresponding model response and evaluation scores.
-MA-SAPO then dynamically retrieves three semantically and thematically relevant exemplars from the training corpus (e.g., “applying ChatGPT in language learning and teaching”).
-These retrieved assets introduce diverse educational contexts and prompt formulations, implicitly guiding refinements such as narrowing task scope and aligning conceptual goals with actionable applications.
-The Analyzer Agent identifies weaknesses in the original prompt notably, vague task specification and overly broad phrasing-both of which risk eliciting generic responses.
-Building on this analysis, the Refiner Agent reformulates the prompt to preserve its educational intent while incorporating a more actionable subtask, such as exploring business opportunities tied to ChatGPT-based assessment practices.
-The optimized prompt exhibits clearer structure, explicit segmentation, and richer contextual depth.
-When regenerated with the refined prompt, the model produces a response that is more coherent, pedagogically grounded, and practically insightful.
-Overall, this case underscores how MA-SAPO integrates multi-agent reasoning and retrieved exemplars to achieve superior prompt optimization, outperforming conventional single-agent or static fine-tuning approaches in terms of coherence, interpretability, and alignment with user intent.
-6. Conclusion
-We propose MA-SAPO, a novel multi-agent framework for interpretable and controllable prompt optimization.
-Compared to prior methods that either perform one-shot rewrites, rely on unguided retrieval, or employ heavy multi-agent debates, MA-SAPO explicitly transforms evaluation outcomes into reusable reasoning assets and grounds refinements in evidence-backed Analyzer to Refiner interactions.
-Our contributions extend beyond framework design to include comprehensive experiments on two benchmarks and multiple backbones, ablation studies on retrieval depth and agent configuration, and human evaluations of reasoning quality and semantic consistency.
-Extensive results demonstrate that MA-SAPO consistently outperforms existing approaches, achieving higher optimization quality while remaining efficient in cost and latency.
-Despite these advantages, MA-SAPO currently depends on the quality of human-annotated scores and a fixed semi-structured schema for asset construction, while also relying on a sparse BM25 retriever and a single reward model-factors that may overlook semantically relevant exemplars and introduce evaluator bias.
-In future work, we plan to improve robustness by integrating a dedicated Feedback Agent to validate asset quality and mitigate bias, as well as by developing hybrid dense–sparse retrieval strategies.
-We also aim to extend the framework to multi-turn prompts to enable more comprehensive evaluation.
+Rather than merely reiterating the optimization pipeline, these examples highlight MA-SAPO’s versatile problem-solving capabilities across different types of prompt deficiencies: (1) structuring broadly scoped queries, (2) contextualizing highly open-ended requests, and (3) disambiguating subjective terminology.
+Across all scenarios, the core multi-agent pipeline remains consistent: specialized agents (e.g., the Analyzer and Refiner) collaboratively evaluate retrieved exemplars to dynamically synthesize a targeted optimization strategy.
+By examining these distinct scenarios, we provide concrete evidence of the framework’s ability to guide language models toward more coherent, interpretable, and intent-aligned responses.
+I.1 Case 1: structuring broadly scoped tasks
+Figure 8: Case 1.
+For a broad query on ChatGPT’s educational impact, retrieved exemplars signal how to frame actionable tasks.
+The system identifies missing operational boundaries and restructures the prompt to demand a specific format (a 10-point list with explanations).
+This structural constraint shifts the final response from generic statements to pedagogically grounded, practical insights.
+I.2 Case 2: contextualizing open-ended requests
+Figure 9: Case 2.
+For an under-specified request for “new novel ideas,” retrieved exemplars highlight the need for a thematic anchor.
+The framework overhauls the prompt by enforcing explicit domain constraints and multi-perspective exploration.
+This enriched context transforms the final response from boilerplate advice into deep, domain-specific conceptualization.
+I.3 Case 3: disambiguating subjective terminology
+Figure 10: Case 3.
+For a query asking for the “strongest” character, retrieved exemplars provide cues for multi-dimensional evaluation.
+The framework refines the prompt by decomposing the subjective term “strongest” into explicit criteria (physical, strategic, and magical abilities).
+These multi-faceted constraints yield a final answer that replaces a simple name-drop with a nuanced, criteria-driven assessment.